### v1

# Inverse sampling intensity weighting for preferential sampling adjustment

Abstract
Traditional geostatistical methods assume independence between observation locations and the spatial process of interest. Violations of this independence assumption are referred to as preferential sampling (PS). Standard methods to address PS rely on estimating complex shared latent variable models and can be difficult to apply in practice. We study the use of inverse sampling intensity weighting (ISIW) for PS adjustment in model-based geostatistics. ISIW is a two-stage approach wherein we estimate the sampling intensity of the observation locations then define intensity-based weights within a weighted likelihood adjustment. Prediction follows by substituting the adjusted parameter estimates within a kriging framework. A primary contribution was to implement ISIW by means of the Vecchia approximation, which provides large computational gains and improvements in predictive accuracy. Interestingly, we found that accurate parameter estimation had little correlation with predictive performance, raising questions about the conditions and parameter choices driving optimal implementation of kriging-based predictors under PS. Our work highlights the potential of ISIW to adjust for PS in an intuitive, fast, and effective manner.
Keywords: model-based geostatistics, preferential sampling, kriging, point process, inverse weighting
1 Introduction
The field of geostatistics comprises a set of methods to infer properties and predict unknown values (or functions) of a spatially continuous process $\{S(\textbf{x}):\textbf{x}\in\mathcal{D}\subseteq\mathbb{R}^{2}\}$ from a discrete set of observation locations, denoted $\textbf{X}=\{\textbf{x}_{1},\textbf{x}_{2},\ldots,\textbf{x}_{n}\}$. Model-based approaches estimated via maximum likelihood estimation (MLE) typically assume $S$ is a realization from a random stochastic process, and subsequent inference and prediction treat X as fixed (Diggle et al.,, 1998). This assumption further implies that the sampled locations are independent of the spatial process of interest, or in symbols, $[\textbf{X},S]=[\textbf{X}][S]$, where square brackets indicate the probability distribution. Violations of this independence assumption are referred to as preferential sampling (PS). There is strong evidence to suggest neglecting PS can negatively impact both geostatistical inference and prediction (Diggle et al.,, 2010; Pati et al.,, 2011; Gelfand et al.,, 2012).
Early works in PS examined X under the lens of a marked point process. Schlather et al., (2004) designed two Monte Carlo tests to detect dependence between marks and their locations. Ho and Stoyan, (2008) introduced the classic PS adjustment model, referred to here as the shared latent process (SLP) model, and derived several of its properties. Full estimation and prediction procedures for the SLP model were formally established in the seminal work by Diggle et al., (2010).
The SLP framework induces dependence between the spatial process of interest and the sampling locations through a shared variable approach. Let $\mu,\alpha,\beta,$ and $\tau^{2}$ be scalars in $\mathbb{R}$, Y denote the $n\times 1$ random vector of observed values sampled at locations X, $S$ the shared latent process, $S(\textbf{X})$ the values of $S$ evaluated at locations X, $IPP$ an inhomogeneous Poisson process, $GP$ a Gaussian process, $C_{\theta}(\cdot,\cdot)$ a stationary covariance function indexed by parameter vector $\theta$, and $\lambda(\cdot)$ the intensity function of X conditional on $S$. Then the SLP model is as follows:
$\begin{split}[\textbf{Y}|\textbf{X},S]\sim N_{n}(\mu\textbf{1}_{n}+S(\textbf{X}),\tau^{2}I_{n}),\\ [\textbf{X}|S]\sim IPP(\lambda),\\ \lambda(\textbf{x})=\exp(\alpha+\beta S(\textbf{x})),\\ S\sim GP(0,C_{\theta}(\cdot,\cdot)).\\ \end{split}$ (1)
While the SLP framework has been extended to handle covariates in the mean model, non-Gaussian likelihoods, and multiple shared latent processes (Pati et al.,, 2011; Watson et al.,, 2019), for the purposes of illustration, we focus on the original formulation of a constant mean, Gaussian likelihood, and one shared latent process. Under the SLP model, sampling locations are no longer assumed to be fixed and are instead a realization from a log Gaussian Cox process (LGCP). The latent process $S$ is assumed to follow a zero mean second order stationary Gaussian process (GP) and for the remainder of this paper, we further assume $C_{\theta}$ to be an isotropic Matérn covariance function. The SLP takes its name from the presence of $S$ in both the mean model of Y and the intensity function of X. The parameter $\beta$ can be interpreted as a single PS coefficient controlling the preferentiality of the SLP. The sign of $\beta$ determines the direction of the preference, while its magnitude determines the likelihood of sampling extreme values.
Subsequent research on PS methods has remained largely faithful to the SLP model. Pati et al., (2011) extended the SLP model structure to a flexible Bayesian framework estimated by MCMC, and proved the posterior consistency of each of the mean, covariance and PS parameters under increasing domain asymptotics. Gelfand et al., (2012) also used Bayesian estimation, and introduced a framework to compare prediction surfaces under PS between different methods. A surprising finding from their simulation analysis was that inclusion of an informative covariate was not sufficient to correct for predictive bias. This empirical finding highlights the importance of the SLP model even when informative covariates explaining the dependence between X and $S$ are available. Dinsdale and Salibian-Barrera, (2019) significantly improved the computational efficiency of the SLP model relative to the previous simulation-based methods by means of the stochastic partial differential equations (SPDE) approach (Lindgren et al.,, 2011) combined with a Laplace approximation implemented in the Template Model Builder (TMB) R package (Kristensen et al.,, 2016). Watson et al., (2019) also fitted the SLP model using the SPDE approach but with an integrated nested Laplace approximation implemented in the R-INLA software (Rue et al.,, 2009), which enjoys comparable computational gains to the approach in Dinsdale and Salibian-Barrera, (2019). The authors further defined a framework to model PS spatio-temporal data and better emulate the evolution of X over time compared to the original SLP model defined in (1) above.
Several other contributions extend the SLP framework to explore a rich set of important applications, including air pollution monitoring (Lee et al.,, 2011, 2015), species distribution modeling (Manceur and Kühn,, 2014; Fithian et al.,, 2015; Conn et al.,, 2017; Pennino et al.,, 2019; Gelfand and Shirota,, 2019; Fandos et al.,, 2021), disease surveillance (Rinaldi et al.,, 2015; Cecconi et al.,, 2016; Conroy et al.,, 2023), phylodynamic inference (Karcher et al.,, 2016), hedonic modelling (Paci et al.,, 2020), bivariate spatial data (Shirota and Gelfand,, 2022), spatially-varying PS (Amaral et al.,, 2023), optimal design under PS (Ferreira and Gamerman,, 2015), space-filling designs (Ferreira,, 2020; Gray and Evangelou,, 2023), a hypothesis test to detect PS in spatio-temporal data (Watson,, 2021), and exact Bayesian inference for the SLP model (Moreira and Gamerman,, 2022; Moreira et al.,, 2023). The improved predictive ability and parameter estimation within the SLP framework contribute to the popularity of model-based approaches accounting for dependence between spatial processes and their observation locations.
While the benefits of the SLP structure and the ubiquity of PS data support its widespread adoption for spatial data analysis, the SLP model still contains significant limitations. Few software packages and out-of-the-box implementations exist for SLP estimation, often coming in the form of custom INLA or MCMC sampler code. Furthermore, the SLP model is challenging to integrate with modern spatial approximation techniques, many of which have become the standard in spatial data analysis due to their scalability and feasibility, offering computationally efficient alternatives to the MLE while maintaining high accuracy (Heaton et al.,, 2019). Among these methods, only the SPDE approach has a well-documented implementation of a PS solution, and incorporating the SLP model into any other method would take a substantial effort. The SLP model requires joint evaluation of both the response and observation likelihood, not a trivial task. Prediction based on the SLP model also can suffer from high computational cost. While parameter estimation for the SLP model may scale well with the SPDE approach, prediction depends on summarizing the posterior distribution of $[S|\textbf{X},\textbf{Y}]$ for each prediction point, which can be infeasible even on a moderately sized grid for both the Laplace approximation and MCMC.
An alternative strategy to the SLP is to incorporate dependence through the sampling intensities $\lambda$ of the observations, rather than the entire likelihood of X. These methods have been developed to much less fanfare compared to the SLP, and it remains unclear how well they work in geostatistical applications. The two main uses of sampling intensities are as a covariate in the mean model of Y or as inverse weights in a likelihood adjustment (Vedensky et al.,, 2023). A potential advantage of sampling intensity methods is their robustness to the form of PS relative to the SLP framework, which addresses a very specific type of clustering process.
The biggest impediment to the use of the sampling intensity is the need to estimate this intensity from the observed locations, the driving reason such approaches have been sidelined in favor of full model-based approaches. Unlike in survey methodology where survey weights are fixed and known, sampling weights in geostatistical applications are largely unknown to the investigator and must be estimated. Unfortunately, nonparametric kernel smoothing estimators of the intensity (Diggle,, 1985; Berman and Diggle,, 1989) have few theoretical guarantees and are not consistent for the true intensity (Guan,, 2008).
Even so, there is preliminary evidence that methods using nonparametrically estimated sampling intensities can still mitigate the effects of PS and, in some scenarios, can predict a surface better than the SLP approach in practice. Reich and Fuentes, (2012) noted how the Bayesian SLP model parameters could be estimated without MCMC by replacing the shared latent process term in the mean model with some function of the sampling intensity $g(\lambda)$. Since the dependence between X and $S$ is completely contained within $g(\lambda)$, the locations X can once again be treated as fixed, so long as $\lambda$ is estimated well. They show improved prediction using the sampling intensity covariate in a 1-dimensional kriging example. In another example, Zidek et al., (2014) provide an approach using estimated weights for air pollution monitoring. Instead of working in a continuous study region, however, these authors considered a finite superpopulation of possible sampling sites and the probability of selecting any site over time was modeled by a logistic regression. The estimated inverse probabilities were then used as weights in a Horvitz-Thompson style design-based estimator for unbiased estimation of parameters under PS. In a third example, Schliep et al., (2023) also considered a finite superpopulation and similarly deviated from traditional model-based geostatistics by using estimated sampling intensity weights to recover the estimated parameter values had the model been fit on the superpopulation, rather than estimate the parameters of the spatial process $S$. These authors also derived a kriging variance estimate adjusted for PS. Finally, Vedensky et al., (2023) conducted a simulation experiment comparing a univariate marginal composite likelihood (CL) weighted by inverse sampling intensities estimated by the MASS::kde2d function in R to the SLP and unweighted CL, and showed improved performance compared to no adjustment. We will refer to methods using inverse sampling intensities as weights in a weighted likelihood adjustment as inverse sampling intensity weighting (ISIW).
Despite preliminary investigations into ISIW for PS adjustment, specifics regarding the effectiveness of such approaches within model-based geostatistics remains largely unknown. In particular, it is unclear whether they reliably estimate both mean and covariance parameters for a latent spatial process on a 2D continuous surface. Composite likelihoods beyond pairwise difference and univariate marginal have also not been explored for ISIW. Finally, the robustness of ISIW and the SLP to misspecification has not been well-studied. In the sections below, we provide an expanded evaluation of ISIW methods, comparing the performance of MLE, SLP, and ISIW applied to the Vecchia and pairwise marginal CLs across multiple random fields and PS designs.
Our work proceeds as follows. We start off in Section 2 by providing background on model-based geostatistics and introduce the key methods we use as a basis for ISIW. In Section 3, we present the implementation of ISIW in detail. In Section 4, we conduct a comparison analysis of the MLE, SLP, and ISIW through a set of simulation studies. In Section 5 we apply the same methods to the famous Galicia moss dataset which has been the dataset of choice when investigating PS. In Section 6 we finish with final remarks and discuss the implications of our work, and outline future directions for continuing investigation.
2 Model-based geostatistics
2.1 Estimation
2.1.1 Maximum likelihood estimation
Under non-preferential sampling (NPS), we define a model for the Gaussian observations Y and underlying Gaussian process $S$ following the SLP framework in (1), but we drop the point process likelihood for X. The standard geostatistical model under NPS follows
$\begin{split}[\textbf{Y}|\textbf{X},S]\sim N_{n}(\mu\textbf{1}_{n}+S(\textbf{X}),\tau^{2}I_{n}),\\ S\sim GP(0,C_{\boldsymbol{\theta}}(\cdot,\cdot)).\\ \end{split}$
We further assume the covariance function $C_{\theta}$ follows a stationary isotropic Matérn covariance defined as
$C_{\theta}(h)=\sigma^{2}\frac{2^{1-\nu}}{\Gamma(\nu)}\left(\sqrt{2\nu}\frac{h}{\phi}\right)^{\nu}K_{\nu}\left(\sqrt{2\nu}\frac{h}{\phi}\right),$
where $\theta:=(\sigma^{2},\nu,\phi)$ defines the covariance parameter vector containing the variance, smoothness, and range parameters respectively, $h$ denotes the Euclidean distance between two points of observation, and $K_{\nu}$ is the modified Bessel function of the second kind. By convention, the smoothness parameter $\nu$ is assumed to be known and fixed a priori before estimation. Therefore, the parameters of interest are $\boldsymbol{\psi}:=(\mu,\sigma^{2},\phi,\tau^{2})$.
Because any finite collection of random variables corresponding to point observations of $S$ follows a multivariate Gaussian distribution and X is treated as fixed, the observed data likelihood is then defined by
$\begin{split}[\textbf{Y},\textbf{X}]\propto N_{n}(\mu\textbf{1}_{n},\Sigma_{n}(\theta)+\tau^{2}I_{n}),\\ \end{split}$ (2)
where $\Sigma_{n}$ is the covariance matrix with $ij$th entry equal to $C_{\theta}(\lVert\textbf{x}_{i}-\textbf{x}_{j}\rVert)$. Estimation proceeds by optimization of the likelihood with respect to the parameter vector $\boldsymbol{\psi}$.
Justification for geostatistical inference via the MLE derives from spatial large sample theory. Two main frameworks have dominated asymptotics for geostatistical estimators: increasing domain asymptotics, and fixed domain (or infill) asymptotics. Increasing domain asymptotics assume that as $n\to\infty$, the study region $\mathcal{D}$ expands, ensuring a minimum separation distance between observation locations and a fixed density of observations per unit area. Fixed domain asymptotics, on the other hand, keep $\mathcal{D}$ fixed while increasing $n$, leading to a growing observation density and vanishing minimum distance between observations (i.e., observations occur closer together as the sample size increases within the fixed study area).
Under increasing domain asymptotics, the MLE for $\boldsymbol{\psi}$ is consistent and asymptotically normal (AN) (Mardia and Marshall,, 1984; Bachoc,, 2014, 2020). Results are more challenging under fixed domain asymptotics due to the inclusion of increasingly close observations of a spatially correlated process. Specifically, for a Matérn Gaussian process with known $\nu$, only a subcomponent of the covariance parameters known as the microergodic parameter is consistently estimable and asymptotically normal (Zhang,, 2004; Kaufman and Shaby,, 2013). We refer to this parameter as $\kappa$, given by $\sigma^{2}/\phi^{2\nu}$ for Matérn covariance functions. The nugget $\tau^{2}$ is also estimable (Tang et al.,, 2021) while the variance $\sigma^{2}$ and range $\phi$ are not (Zhang,, 2004). Although the smoothness parameter $\nu$ is theoretically estimable, it is numerically challenging to estimate and is conventionally treated as fixed (Loh et al.,, 2021). Fixed effect coefficients in the mean are typically non-estimable (Wang et al.,, 2020), except under specific smoothness conditions on the covariates (Yu,, 2022; Bolin and Wallin,, 2024; Gilbert et al.,, 2024).
The main drawback for practical use of the MLE is its computational burden. Evaluating the likelihood requires an inversion of the covariance matrix, which has $O(n^{3})$ time and $O(n^{2})$ space complexity making it infeasible for applications for moderately sized $n$. Recent advancement in modern spatial statistics has focused on GP parameter estimation using approximations which massively reduce computational burden in exchange for marginal decreases in efficiency. We explore composite likelihood and the Vecchia approximation, two approaches which have been widely adopted in geostatistics and are particularly conducive to weighting.
2.1.2 Composite likelihood
Sticking with our previous notation, let $\textbf{Y}=(Y_{1},Y_{2},...,Y_{n})^{\top}\in\mathbb{R}^{n}$ be a $n\times 1$ random vector with probability density $f(\textbf{y};\boldsymbol{\psi})$ for unknown $r$-dimensional parameter vector $\boldsymbol{\psi}\in\Psi\subseteq\mathbb{R}^{r}$. Define $\{\mathcal{A}_{1},...,\mathcal{A}_{K}\}$ to be a set of $K$ marginal or conditional events with associated likelihood $\mathcal{L}_{k}(\theta;\textbf{y})\propto f(\textbf{y}\in\mathcal{A}_{k};\boldsymbol{\psi})$. The log composite likelihood (CL) is the weighted sum of the event-specific log likelihoods
$\log\mathcal{L}_{C}(\boldsymbol{\psi};\textbf{y}):=\sum_{k=1}^{K}w_{k}\log\mathcal{L}_{k}(\boldsymbol{\psi};\textbf{y}),$
where $\textbf{w}:=(w_{1},...,w_{K})$ is a vector of weights, not necessarily non-negative and the density $f$ follows that defined in (2). The maximum CL estimate is defined as $\hat{\boldsymbol{\psi}}_{C}:=\text{argmax}_{\boldsymbol{\psi}}\log\mathcal{L}_{C}(\boldsymbol{\psi};\textbf{y})$. One common choice of CL is the univariate marginal where the log likelihood is the sum of the log marginal density of each observation.
$\log\mathcal{L}_{UM}(\boldsymbol{\psi};\textbf{y})=\sum_{i=1}^{n}w_{i}\log f(y_{i};\boldsymbol{\psi}).$
Typically, the use of this univariate marginal in geostatistics is limited because it ignores dependencies between observations and thus cannot estimate covariance parameters. For this reason, pairwise CLs have been much more popular (Varin et al.,, 2011). Bevilacqua and Gaetan, (2015) compared the efficiency of three different pairwise CL estimators: pairwise marginal (PMLE), pairwise conditional (PCMLE), and pairwise difference (PDMLE) CLs. These authors concluded that the PMLE outperformed all other pairwise CLs and recommended its use over the PCMLE and PDMLE. The likelihoods for each pairwise CL are
$\begin{split}\log\mathcal{L}_{PM}(\boldsymbol{\psi};\textbf{y})=\sum_{i<j}w_{ij}\log f(y_{i},y_{j};\boldsymbol{\psi}),\\ \log\mathcal{L}_{PC}(\boldsymbol{\psi};\textbf{y})=\sum_{i\neq j}w_{ij}\log f(y_{i}|y_{j};\boldsymbol{\psi}),\\ \log\mathcal{L}_{PD}(\boldsymbol{\psi};\textbf{y})=\sum_{i<j}w_{ij}\log f(y_{i}-y_{j};\boldsymbol{\psi}).\end{split}$ (3)
These authors also proved, under increasing domain asymptotics, the consistency and asymptotic normality of $\boldsymbol{\hat{\psi}}_{C}$ estimated from (3). Bachoc et al., (2019) proved the same properties for the PMLE and PCMLE in the one-dimensional setting under fixed domain asymptotics. However, the true utility of CL lies in its computational speed and robustness to misspecification. While the MLE requires a $O(n^{3})$ matrix inversion and specification of the joint density, pairwise CL consists only of $O(n^{2})$ terms and only requires correct specification of second order densities. The computational efficiency of pairwise CL can be further enhanced by using the weights w to only include pairwise observations within a certain distance $d$ apart (Bevilacqua and Gaetan,, 2015).
In later sections, we will use the PMLE as a basis for ISIW due to its superior performance over other pairwise likelihoods. In particular, we do not consider the univariate marginal or PDMLE because they do not directly estimate the entirety of $\boldsymbol{\psi}$. In the next section, we discuss the Vecchia approximation, an alternative likelihood approach that achieves even greater computational and statistical efficiency than the CL.
2.1.3 Vecchia approximation
The Vecchia approximation is a specific case of CL based on the observation that the joint distribution can be decomposed as the product of conditional distributions. Let $\textbf{Y}=(Y_{1},...,Y_{n})^{\top}\in\mathbb{R}^{n}$ and $p:\{1,...,n\}\to\{1,...,n\}$ be a permutation mapping which reorders Y. We define the history of variable $Y_{j}$ as a random subvector $\textbf{Y}_{h(j)}$ where $h(j)=\{l\in\{1,...,n\}:p(l)<p(j)\}$. The joint density of Y can then be refactored as
$f(\textbf{y};\boldsymbol{\psi})=f(y_{p(1)};\boldsymbol{\psi})\prod_{i=2}^{n}f(y_{p(i)}|\textbf{y}_{h(i)};\boldsymbol{\psi}).$
Vecchia observed that much of the information in the conditioning sets with higher indices was likely to be redundant. One could attain a good tradeoff of efficiency for computational gain by decreasing the size of each conditioning set and being judicious about which variables to include. The density would then be approximated as
$f(\textbf{y};\boldsymbol{\psi})\approx{f}_{V}(\textbf{y};\boldsymbol{\psi})=f(y_{p(1)};\boldsymbol{\psi})\prod_{i=2}^{n}f(y_{p(i)}|\textbf{y}_{q(i)};\boldsymbol{\psi}),$ (4)
where $q(i)\subseteq\{p(1),p(2),...,p(i-1)\}$ is the set of indices constituting the conditioning set of $y_{p(i)}$. The Vecchia estimate $\hat{\boldsymbol{\psi}}_{V}$ maximizes (4). The Vecchia approximation can also be written in a weighted CL form,
$\begin{split}\log\mathcal{L}_{V}(\boldsymbol{\psi};\textbf{y})&=w_{1}\log f(y_{p(1)};\boldsymbol{\psi})\\ &+\sum_{i=2}^{n}w_{1i}f(y_{p(i)},\textbf{y}_{q(i)};\boldsymbol{\psi})\\ &-\sum_{i=2}^{n}w_{2i}\log f(\textbf{y}_{q(i)};\boldsymbol{\psi}).\end{split}$ (5)
While simple in concept, the Vecchia approximation requires careful selection of three key hyperparameters: 1) the size of the conditioning set, 2) which variables to include in each conditioning set, and 3) the ordering of the variables as determined by $p$. We denote the maximum size of any $q(i)$ as $m$. In all later sections, we choose $m=20$ based on a clear case of diminishing returns in inference and prediction for $m>20$ observed empirically in Datta et al., (2016). Default settings in two implementations of the Vecchia approximation (GPvecchia and GpGp R packages) are $m=20$ and $m=30$, respectively (Katzfuss and Guinness,, 2021; Guinness,, 2021). For the choice of which variables to include in the neighborhood $q(i)$, we follow the recommendation given in Vecchia, (1988) by choosing the nearest neighbors to $y_{p(i)}$ measured by Euclidean distance. Ordering for the Vecchia approximation requires additional specification in multidimensional settings due to the lack of natural ordering in observations of a spatial point process. We use the maxmin ordering scheme, which picks a location $p(i)$ sequentially by maximizing the distance to the nearest point in $\{y_{p(1)},y_{p(2)},...,y_{p(i-1)}\}$ and has been shown to approximate the true distribution better than other ordering methods (Guinness,, 2018; Katzfuss and Guinness,, 2021).
Theoretical justification of the Vecchia approximation in spatial statistics is relatively light compared to that of the MLE and CL. Typical approaches to proving fixed domain asymptotics fail due to the loss of stationarity in the parent process introduced by the Vecchia approximation (Zhang et al.,, 2024). However, it is impossible to deny its excellent empirical performance in applications (Heaton et al.,, 2019). A key advantage the Vecchia approximation possesses over typical CL is that (4) and (5) correspond to a valid joint probability distribution for Y. As a result, the Kullback-Leibler (KL) divergence of the Vecchia approximation with respect to the true distribution can be computed and has been shown to be a nonincreasing function of $m$ when the conditioning sets $q(i)$ are chosen as the nearest neighbors (Katzfuss et al.,, 2020; Huser et al.,, 2023).
The Vecchia approximation achieves a good balance of computational complexity and efficiency. It clearly outperforms other CL methods in terms of what order correlations are being accounted for, scalability, and probabilistic validity. This is the main reason for considering its use in ISIW methods compared to CL, a point we expand on in later sections.
2.2 Prediction
For many spatial analyses, the main goal is to predict values at unobserved locations. We focus on spatial prediction via kriging, or Gaussian process regression, a function approximation technique with point and variance estimate defined as
$\hat{S}(\textbf{X}_{0})$ $=\mu+C_{\theta}(\textbf{X}_{0},\textbf{X}_{n})^{\top}\boldsymbol{\Sigma}_{n}(\theta)^{-1}(\textbf{Y}-\mu\textbf{1}_{n}),$ (6)
$\operatorname{Var}(\hat{S}(\textbf{X}_{0}))$ $=C_{\theta}(\textbf{X}_{0},\textbf{X}_{0})-C_{\theta}(\textbf{X}_{0},\textbf{X}_{n})^{\top}\boldsymbol{\Sigma}_{n}^{-1}(\theta)C_{\theta}(\textbf{X}_{0},\textbf{X}_{n}),$
where $\textbf{X}_{0}$ represents unobserved locations to be predicted. In practice, since the parameters are unknown, all parameters in (6) are replaced with estimates obtained from any of the aforementioned or other estimation methods.
Kriging has several properties that make it an effective predictor of spatial surfaces. As the best linear unbiased predictor (BLUP), it minimizes mean squared prediction error among all unbiased predictors (Cressie and Wikle,, 2015). However, it is important to note that this optimality no longer holds under preferential sampling, where even the true values of $\boldsymbol{\psi}$ used in the kriging equations yield suboptimal predictions (Dinsdale and Salibian-Barrera,, 2019). Stein established the asymptotic efficiency of kriging under various robustness conditions (Stein,, 1988; Stein, 1990a, ; Stein, 1990b, ; Stein,, 1993), while Wang et al., (2020) showed that kriging’s prediction error vanishes under a uniform metric. Additionally, Putter and Young, (2001) demonstrated that the difference between predictions using kriging with estimated and the true parameters is asymptotically negligible if the joint Gaussian distributions of the spatial process under the true and estimated covariance functions are contiguous almost surely. Given these results, one could argue that the geostatistical model in (2) is more appropriate for spatial interpolation rather than inference.
3 Inverse sampling intensity weighting
We next describe estimation and prediction procedures of ISIW for PS adjustment. ISIW is a two-stage approach, wherein we first estimate the sampling intensity at each of the observation locations X, and, in the second stage, we input the (estimated) inverse sampling intensities as weights into a weighted likelihood adjustment. The resulting adjusted parameter estimates are then substituted in the kriging equations for prediction.
3.1 Estimation of sampling intensity
The essence of ISIW is to account for the dependence between the response and observation locations by using the vector of estimated sampling intensities instead of the more complex full likelihood of the observation process. This approach assumes that the locations contain sufficient information about the spatial dependence structure of the observation locations to enable proper adjustment.
Table 1: Point process intensity function estimators considered in the simulation study.
Method Reference Bandwidth Selection
diggle Diggle, (1985) Least-squares cross-validation
scott Scott, (1992) Rule-of-thumb based on normal reference density
ppl Loader, (1999) Likelihood cross-validation (leave-one-out)
CvL Cronie and Van Lieshout, (2018) Maximum likelihood cross-validation
CvL.adaptive van Lieshout, (2021) Adaptive bandwidth based on local point density
R-INLA Simpson et al., (2016) LGCP model (non-kernel-based)
Methods to estimate a spatially varying intensity of a point process can be divided into parametric and nonparametric approaches. Domain knowledge can inform the parametric form of the intensity either through choice of model or covariates, but oftentimes this information is unavailable and nonparametric estimation is required. The nonparametric approaches follow the kernel smoothing approach discussed in Diggle, (1985). Let $k$ be a $d$-dimensional kernel function from $\mathbb{R}^{d}\to\mathbb{R}^{+}$, which is a symmetric probability density function. Given a bandwidth size $h>0$ and edge correction factor $w_{h}(\cdot,\cdot)$, the nonparametric kernel smoothing estimator of the intensity function is given by
$\hat{\lambda}(\textbf{x};h)=h^{-d}\sum_{\textbf{y}\in\textbf{X}\cap\mathcal{D}}k\left(\frac{\lVert\textbf{x}-\textbf{s}\rVert}{h}\right)w_{h}(\textbf{x},\textbf{s})^{-1},\textbf{x}\in\mathcal{D}.$
The key hyperparameter for nonparametric intensity estimation is the bandwidth size $h$. Bandwidth selection is a well-studied problem, with methods ranging from high bandwidth, smooth intensity estimators to low bandwidth, flexible intensity estimators. Each has its place in the bias-variance tradeoff, with higher bandwidths exhibiting higher bias with lower variance and lower bandwidths vice versa. We experiment with several bandwidth selection strategies implemented in the spatstat R package (Table 1).
The effectiveness of any ISIW approach for PS should improve the closer estimated weights $\lambda(\textbf{x})^{-1}$ are to the true inverse intensities. In theory, a parametric LGCP model should estimate the true intensity better than any nonparametric estimator if X follows a LGCP, especially with multiple realizations. However, it is known that it is impossible to distinguish between a single realization from an inhomogenous Poisson process with deterministic intensity $\lambda(\textbf{x})$ and one from a stationary Cox process whose conditional intensity coincides with $\lambda$ (Gelfand et al.,, 2010). We consider a LGCP method introduced in Simpson et al., (2016) estimated by R-INLA to verify whether a known parametric model shows improvement over a nonparametric estimator.
Once the sampling intensities are estimated at every observation location, we calculate weights as the inverse of the sampling intensities and then normalize these to sum up to the number of observations $n$. The estimated weight for the $i$th observation becomes
$\hat{w}_{i}=\hat{w}(\textbf{x}_{i})=n*\frac{\hat{\lambda}(\textbf{x}_{i})^{-1}}{\sum_{\textbf{s}\in\textbf{X}}\hat{\lambda}(\textbf{s})^{-1}}.$ (7)
3.2 Defining the likelihood
ISIW applies weights proportional to the inverse sampling intensity to each event in a likelihood factored as a product of densities. The only known examples of ISIW in the PS literature include a pairwise difference (Schliep et al.,, 2023) and univariate marginal CL (Vedensky et al.,, 2023). The pairwise difference CL cannot estimate $\mu$ directly, limiting ISIW’s impact on PS adjustment of the mean, while the univariate marginal CL cannot estimate covariance parameters $\theta$. To address these gaps, we propose two new ISIW estimators for simultaneous mean and covariance estimation: the ISIW pairwise marginal (ISIW-PM) and ISIW Vecchia (ISIW-V) estimator. Let the weights be defined as in (7). Then the ISIW-PM follows
$\begin{split}\log\mathcal{L}_{WPM}(\boldsymbol{\psi};\textbf{y})=\sum_{i<j}w_{i}w_{j}\log f(y_{i},y_{j};\boldsymbol{\psi}).\end{split}$
For the ISIW-V, we initially considered weighting the likelihood as follows
$\begin{split}\log\mathcal{L}_{WV}(\boldsymbol{\psi};\textbf{y})&=w_{p(1)}\log f(y_{p(1)};\boldsymbol{\psi})\\ &+\sum_{i=2}^{n}\left(\prod_{j\in\{p(i)\}\cup q(i)}w_{j}\right)\log f(y_{p(i)},\textbf{y}_{q(i)};\boldsymbol{\psi})\\ &-\sum_{i=2}^{n}\left(\prod_{j\in q(i)}w_{j}\right)\log f(\textbf{y}_{q(i)};\boldsymbol{\psi}),\end{split}$
to maintain the probabilistic interpretation of the inverse weighting. However, numerical issues arise due to taking the product of several small weights, greatly increasing the chance of extreme values. As a more computationally efficient approach, we approximate the true weight using the following weighted Vecchia approximation
$\begin{split}\log\mathcal{L}_{WV}(\boldsymbol{\psi};\textbf{y})&=w_{p(1)}\log f(y_{p(1)};\boldsymbol{\psi})\\ &+\sum_{i=2}^{n}w_{p(i)}\log f(y_{p(i)},\textbf{y}_{q(i)};\boldsymbol{\psi})\\ &-\sum_{i=2}^{n}w_{p(i)}\log f(\textbf{y}_{q(i)};\boldsymbol{\psi}).\end{split}$
3.3 Numerical estimation
The ISIW for both the Vecchia and composite likelihood can be estimated by standard optimization procedures. In our study, we used the L-BFGS-B routine as implemented in the optim package in the R language. Initial values for parameters to be estimated were selected with general rules of thumb from the GPVecchia R package and all other optimization parameters were set to their default settings.
3.4 Winsorization of extreme weights
As with other inverse weighting procedures, a key concern for ISIW is extreme weights. Observation locations in sparsely sampled areas will have extremely high weights in the final likelihood and can cause numerical issues in the optimization. Two straightforward options are trimming and Winsorization. Trimming involves eliminating values considered as outliers, while Winsorizing enforces outliers to a threshold value. We chose Winsorization to handle extreme values because trimming observations can remove rare information not captured by less extreme weights, weakening the effectiveness of PS adjustment. For our Winsorization procedure, we first normalize the sampling intensities $\hat{\lambda}$ to sum to $n$. We then Winsorized the normalized intensities by choosing a lower threshold (to be elaborated on later), and normalized the new resultant weights to sum to $n$.
3.5 Prediction
[…]
4 Simulation analysis
4.1 Experiment
We conducted a simulation experiment to evaluate the inferential and predictive performance of ISIW-PM and ISIW-V versus the benchmark MLE and SLP approaches under preferential sampling. We generated $B=500$ realizations for two separate Matérn random fields on a $200\times 200$ grid on the unit square: a low-range ($\phi=0.02$) and high-range ($\phi=0.15$) surface with $\mu=4,\sigma^{2}=1.5,\nu=1,$ and $\tau^{2}=0.1$. To test robustness of methods against a variety of observation patterns, we sampled three different point processes conditional on having $n=100$ and $n=800$ observations: a log Gaussian Cox process (LGCP), a sigmoidal Cox process (SCP), and a Thomas process. Traditionally, the SLP approach has only been tested against LGCP observation patterns where it is correctly specified.
The intensity function for the LGCP is $\exp\{\beta S(\mathbf{x})\}$. For the sigmoidal Cox process, the intensity function follows $\beta(1+\exp\{-S(\mathbf{x})\})^{-1}$. For the Thomas process, a set of parent points was first generated from a homogeneous Poisson process. The number of offspring for each parent then followed a Poisson distribution with mean $\exp\{\beta S(\mathbf{x})\}$ and offspring points were sampled from a normal distribution centered at their respective parent locations, with a scale parameter of 0.1. We set $\beta=1$ for all point processes. The SCP, LGCP, and Thomas process were deliberately selected to represent visually similar clustered point processes, ordered by increasing clustering intensity, with more distinct and separated clusters in the Thomas process (Figure 1).
Figure 1: One realization of the point processes used in the simulation analysis for $n=200$. The top and bottom row fields are low range ($\phi=0.02$) and high range ($\phi=0.15$) respectively.
For each set of observations, we fit the MLE, SLP, ISIW-PM, and ISIW-V models and compared the results in terms of parameter estimation and prediction. The MLE, ISIW-PM, and ISIW-V were estimated using optim as described in Section 3.3. We estimated the SLP model parameters using the INLA-SPDE implementation as defined in Watson et al., (2019) and used penalized complexity (PC) priors on the variance and range parameters (Simpson et al.,, 2017; Fuglstad et al.,, 2019). For the range parameter we set $P(\phi<0.01)=0.01$ and for the variance parameter we set $P(\sigma>10)=0.01$. The default priors for INLA were used for $\mu$ and $\tau$, set as $N(0,1000)$ and $\text{Gamma}(1,5\times 10^{-5})$, respectively. Prediction for the MLE and all ISIW methods followed the kriging equations, with point estimates of the requisite parameters substituted into (6).
We evaluated several variants of the ISIW methods, distinguished by their first-stage sampling intensity weight estimation (Table 1). As a gold standard for assessing the optimal performance of ISIW, we also used the true weights extracted from the simulation surface, referred to as “Known” weights. The intensity estimators included multiple nonparametric kernel density estimators from spatstat and a parametric LGCP model, which used the same priors and mesh as the SLP. To determine a threshold for Winsorization, we conducted a small numerical experiment, simulating a Matérn random field with high range ($\phi=0.15$) 100 times. For each replicate, we generated two LGCP point patterns with $\beta=1$, one with $N=100$ and the other with $N=800$. ISIW-V was then fit under different weight thresholds using the bw.diggle bandwidth selection, as it was the most prone to numerical instability. The best-performing threshold, $1\times 10^{-2}$, was selected based on yielding the best average root mean-squared prediction error (RMSPE). All subsequent ISIW methods implemented Winsorization using the $1\times 10^{-2}$ threshold.
Our approach to evaluating each method prioritized predictive performance due to the aforementioned challenges with standard statistical inference in fixed domain asymptotics and preferential sampling. The RMSPE was computed over each center point in a $200\times 200$ grid, yielding 40,000 prediction points per method and simulation scenario. Parameter estimation was then assessed by computing relative bias and relative root mean-squared error (RMSE) for the parameter vector $(\mu,\sigma^{2},\phi,\tau^{2},\kappa)$.
4.2 Results
Across all simulation scenarios, INLA-SLP and ISIW-V Known were the top two predictors by RMSPE, with similar performance to each other and a clear margin over other methods. In general, INLA-SLP had better predictive performance in the high range scenario, whereas ISIW-V Known did so in the low range setting (Tables S1 and S2). The one exception was the high range scenario for the $N=100$ Thomas process, where ISIW-V Known had both a lower RMSPE (0.694) and a lower standard deviation (0.15) compared to INLA-SLP (0.749 and 0.27) (Table S3). However, there were instances of predictive instability in INLA-SLP, as indicated by the high standard deviation of RMSPE, particularly in the $N=100$ low range SCP and high range Thomas process $N=100$ scenarios where INLA-SLP was misspecified. The ISIW methods also tended to run faster than the INLA-SLP at both sample sizes and had similar runtimes to the MLE (Table S4). We did not experiment with larger sample sizes than $N=800$ because PS is primarily a concern in small to moderate sample sizes; as sample sizes grow sufficiently large, its effects diminish.
Table 2: Predictive performance of all sixteen methods based on median rank, mean rank, and percentage of total simulations when RMSPE for the method was lower than that of MLE and SLP. Rank was determined by RMSPE.
Method Median Rank Mean Rank % RMSPE lower % lower RMSPE
of RMSPE of RMSPE than MLE than SLP
INLA-SLP 2 5.29 73.3 NA
ISIW-V Known 2 5.75 70.0 44.0
ISIW-V CvL.adaptive 5 6.53 72.7 31.8
ISIW-V diggle 5 5.96 78.5 29.3
ISIW-V CvL 7 7.38 85.0 28.9
ISIW-V ppl 7 7.37 74.4 27.8
ISIW-V INLA 8 8.23 74.0 27.9
ISIW-PM CvL.adaptive 8 8.01 65.1 28.3
ISIW-PM diggle 8 7.98 61.3 28.1
ISIW-V scott 9 8.29 85.6 28.4
ISIW-PM Known 9 8.60 51.8 23.9
ISIW-PM ppl 10 9.84 50.4 27.9
MLE 11 10.2 NA 26.7
ISIW-PM CvL 12 11.3 30.9 26.5
ISIW-PM INLA 14 12.3 27.0 24.9
ISIW-PM scott 14 12.7 22.3 24.6
While ISIW-V using estimated weights did not match the predictive performance of ISIW-V Known, weight estimation variants including diggle and CvL.adaptive showed improved predictive performance over the alternative of no PS adjustment (Table 2). They achieved a median rank of 5 out of all 16 methods, had a lower RMSPE than MLE in over 70% of simulations, and even outperformed SLP nearly 30% of the time (for comparison, ISIW-V Known outperformed SLP in 44% of simulations). The CvL variant outperformed MLE in 85% of simulations, surpassing even SLP, which did so in 73%. Surprisingly, ISIW methods using weights estimated by the parametric LGCP model fitted in INLA performed relatively poorly, with a median rank of 8, though they still consistently outperformed MLE. All ISIW-V variants outperformed MLE in over 70% of simulations, whereas ISIW-PM lagged far behind, with some variants on average ranking worse than the MLE.
The variants that predicted best tended to estimate lower bandwidths which introduced greater noise in prediction, as reflected in the higher standard deviation of RMSPE for these variants. In contrast, scott and ppl traded variance for bias, which increased numerical stability but decreased the impact of weighting, making them nearly indistinguishable from prediction based on the MLE.
Figure 2: Example predictive surface for the MLE, INLA-SLP, and ISIW-V approaches. The first panel shows the true field values with observations denoted as points, while the remaining panels display residuals (observed minus predicted values) for each method.
Figure 2 provides an illustrative example comparing the predictive tendencies of MLE, INLA-SLP, and ISIW-V Known. In the data-rich area at the bottom, predictions from all three methods are nearly identical. The differences emerge in data-sparse regions where PS is known to have its greatest impact. While the MLE overpredicts the center of the grid, the INLA-SLP predicts a much lower value over the same region which decreases predictive error but leads to underestimation in a specific area indicated by the deep purple. ISIW-V, similar to INLA-SLP, lowers predictions in that region but not as drastically, avoiding the central underprediction seen in INLA-SLP but still overestimating in another area, shown as dark orange slightly to the right of the center. As shown by this example, both INLA-SLP and ISIW-V adjust for PS in similar ways, but INLA-SLP tends to apply a larger adjustment in data sparse areas compared to ISIW-V.
Table 3: Relative bias and RMSE in parameter estimation for MLE, SLP, and ISIW methods across all simulation scenarios. Bolded values indicate the method with the smallest bias or RMSE for a given parameter.
Method Variant $\mu$ $\sigma^{2}$ $\phi$ $\tau^{2}$ $\kappa$
Bias RMSE Bias RMSE Bias RMSE Bias RMSE Bias RMSE
MLE - 0.106 0.173 0.046 0.589 0.013 0.429 -0.057 0.897 0.113 0.538
INLA-SLP - 0.016 0.116 0.241 1.32 0.110 0.434 $2.13\times 10^{2}$ $1.19\times 10^{4}$ 0.546 17.1
ISIW-V Known -0.259 0.409 0.035 20.8 90.9 $4.52\times 10^{3}$ 0.053 1.72 0.712 1.58
CvL 0.022 0.240 -0.057 1.10 65.1 $2.71\times 10^{3}$0.007 1.14 0.140 0.624
CvL.adaptive -0.018 0.268 0.362 51.9 $1.25\times 10^{3}$ $3.08\times 10^{4}$ 0.441 2.44 2.44 1.14
INLA 0.059 0.170 -0.068 0.807 -0.023 0.673 -0.013 1.15 0.166 0.804
ISIW-PM Known -0.161 0.309 -0.293 0.408 -0.352 0.489 0.713 1.09 1.900 7.65
CvL 0.122 0.203 0.061 0.596 -0.168 0.397 0.871 1.50 0.885 2.05
CvL.adaptive 0.042 0.150 -0.162 0.511 -0.184 0.630 0.782 1.35 1.550 11.0
INLA 0.135 0.211 -0.018 0.570 -0.201 0.414 0.859 1.56 0.912 1.69
Table 3 compares the relative bias and relative RMSE in parameter estimation for the MLE, SLP, and ISIW methods, with a specific focus on the Known, CvL, CvL.adaptive, and INLA variants. While the MLE approach overestimated $\mu$ and INLA-SLP estimated $\mu$ accurately, ISIW-V Known significantly underestimated the mean by over 25% (averaging an estimate of three for a true mean of four), with a highly variable RMSE of 0.409. The estimated weight variants of ISIW-V had higher error in mean estimation compared to INLA-SLP but lower error than ISIW-V Known and MLE.
For the range parameter, ISIW-V estimates were consistently too high across all variants and exhibited potential numerical instability. In contrast, ISIW-PM underestimated the range but did so with much greater stability. MLE had relatively low estimation error for all covariance parameters, particularly $\kappa$ and $\tau^{2}$, aligning with theoretical expectations. Aside from the mean, INLA-SLP did not estimate any parameter particularly well, with $\tau^{2}$ showing significant numerical instability.
There was no clear relationship in the simulation between the method of parameter estimation and resulting predictive performance. INLA-SLP estimated the nugget and microergodic parameter extremely poorly, despite these being theoretically among the easiest to estimate, while ISIW-V on average underestimated the mean by 25% and vastly overestimated the range by more than a factor of 90. Yet, both methods led in predictive performance. In contrast, MLE and the CvL and INLA variants of ISIW produced relatively good parameter estimates but consistently ranked lower in predictive performance, with prediction errors far exceeding those of INLA-SLP and ISIW-V.
5 Application to the Galicia moss data
Figure 3: Observed lead concentrations at sampled locations in Galicia in 1997 and 2000.
The Galicia moss dataset has been a widely used example for illustrating the effects of preferential sampling on geostatistical inference and prediction (Diggle et al.,, 2010; Dinsdale and Salibian-Barrera,, 2019). It contains lead concentrations in moss samples, measured in micrograms per gram of dry weight, collected from Galicia, northern Spain. Figure 3 shows the spatial distribution of 63 measurements taken in 1997 and 132 measurements taken in 2000. Sampling in 1997 was preferential, with a bias toward locations in the north with lower lead concentrations, whereas sampling in 2000 was more regular and non-preferential. Further details on the dataset can be found in Diggle et al., (2010).
We fit each of the MLE, INLA-SLP, ISIW-V, and ISIW-PM approaches using the 1997 and 2000 data separately and generated predictive surfaces over a $20\times 20$ grid covering the Galicia area. Based on its good performance in the simulation analysis, we estimated the ISIW first stage intensities using CvL.adaptive bandwidth selection. We did encounter numerical difficulties when fitting the INLA-SPDE model to the 1997 data. While previous analyses of the Galicia dataset have used areal models, we elected to use the SPDE approach to be consistent with our simulation analysis. The model was particularly sensitive to the choice of prior for the range and initial value for the nugget. It would sometimes generate predictive surfaces with no variation, where the same value was predicted across the entire region including in locations with observations recording different values. Here we present the predictive surface that exhibited some variation and aligned with expectations from kriging, where predictions have similar values to the observations closeby.
Figure 4: Spatial predictions of lead concentrations in Galicia in 1997 and 2000 using MLE, INLA-SLP, ISIW-V, and ISIW-PM. Points represent the observed data.
Figure 4 displays the predictions for the lead concentration surface. There was little difference for the 2000 data, however, substantial differences emerged in the 1997 data. As expected, both the ISIW-V and INLA-SLP estimated higher values in data sparse areas to correct for the preferential sampling of low values. The INLA-SLP applied a more extreme correction than the ISIW-V, similar to the example in Figure 2. It unexpectedly predicted values exceeding 10.5 in the south, where the nearest observations fall below 5.00. It also estimated a very high $\mu$ of 10.79 (Table S5). This was unusual considering the maximum observation was 9.51 and the mean of all 63 observations in 1997 was 4.72. The MLE, ISIW-V, and ISIW-PM only estimated a $\mu$ of 5.15, 6.07, and 5.95, respectively. Besides being a result of INLA-SLP’s PS adjustment, the discrepancy could also be attributed to numerical instability in INLA-SLP estimation, and other priors or initial values might produce different results.
Similar to our simulation results, the ISIW-PM estimated a lower range compared to other methods. This low range likely explains why its predictive surface exhibits little spatial dependence with nearly uniform predictions across the region. ISIW-V appears to strike a balance between implementing a moderate adjustment for PS while preserving the spatial correlation evident in the dataset.
6 Discussion
In this work, we compared ISIW methods to the SLP and MLE for parameter estimation and prediction under preferential sampling. We implemented ISIW in a novel combination with the Vecchia approximation and found that, despite being less well-suited for geostatistical inference, ISIW with the Vecchia approximation nevertheless improves predictive accuracy over the MLE approach and remains competitive with the INLA-SLP approach using the true weights. We conducted a comprehensive simulation study that extended beyond previous analyses, evaluating the robustness of ISIW across various settings. Its benefits persisted across multiple observation point processes, different levels of spatial dependence in the underlying field, and alternative point process intensity estimators used in the first stage of ISIW. We also showed an application to the Galicia moss data where ISIW appears to give more reasonable predictions compared to the standard SLP approach.
Several factors make ISIW an attractive alternative to SLP for PS adjustment, namely its speed, simplicity, and strong predictive performance. Not only does it scale more efficiently with increasing $n$, but in terms of the practical implementation, we found that our version of ISIW-V ran faster than INLA-SLP for combined estimation and prediction, even though we did not implement an optimized version. ISIW is also simple to use and understand. INLA and TMB require specification of both the prior distributions for hyperparameters and mesh triangulation that can significantly alter results (Righetto et al.,, 2020). In contrast, ISIW is a straightforward inverse weighting method familiar to much of the research community that only requires an estimated vector of weights. Furthermore, since most effective Gaussian process approximation techniques are already based on the Vecchia approximation (Heaton et al.,, 2019), integrating ISIW into state-of-the-art methods is significantly simpler and more straightforward compared to the SLP. These computational and conceptual advantages come without sacrificing predictive performance. In our simulation analysis, ISIW achieved improved predictive accuracy over the MLE with estimated weights and performed comparably to the gold standard SLP when the true weights were known. In scenarios where the SLP was misspecified, ISIW outperformed it, demonstrating greater robustness and fewer parametric assumptions.
Many improvements can still be made to ISIW. A significant gap remains between the predictive accuracy achieved by ISIW with estimated and true weights. The lack of accuracy and theoretical guarantees for the nonparametric intensity estimators used in this study may contribute to their inability to achieve similar performance. Moreover, Winsorization is necessary to prevent pathological numerical behavior but remains an undesirable ad hoc step. Intensity modelling incorporating the values of the observations rather than just their locations may also lead to improved weight estimation. Future work should explore more flexible and robust estimators for point process intensities beyond kernel density estimators and the INLA approach.
Additionally, ISIW does not yield reliable parameter estimates for the underlying geostatistical model. The weighted likelihood adjustment introduces substantial bias and error into parameter estimation, yet unexpectedly improves predictions through kriging. The apparent disconnect between parameter estimation and prediction warrants further investigation. It remains unclear why parameters obtained from the weighted likelihood adjustment consistently lead to better kriging predictions. While kriging is known to lose its predictive optimality under preferential sampling, the results of this study provide insight into conditions under which it can still serve as an effective function approximation tool under relaxed assumptions like PS.
Finally, we did not address uncertainty quantification for parameter estimates or predictions. While the kriging variance formula remains available and could be computed using the estimated parameters, its utility in the presence of preferential sampling is uncertain. In addition, geostatistical inference is already challenging and preferential sampling adjustment appears to exacerbate the difficulty. As a result, any standard errors for parameter estimates obtained under ISIW are unlikely to have valid coverage properties. Future work will focus on developing principled approaches to quantifying uncertainty in both inference and prediction under ISIW.
7 Supplementary Material
Table S1: Average RMSPE (standard deviation) for methods under the LGCP simulation. Bolded values indicate the method with the lowest average RMSPE across all simulations for that setting.
N = 100 N = 800
Method Weights $\phi=0.02$ $\phi=0.15$ $\phi=0.02$ $\phi=0.15$
MLE 1.726 (0.11) 0.785 (0.17) 1.306 (0.06) 0.361 (0.07)
INLA-SLP 1.195 (0.08) 0.553 (0.09) 0.995 (0.04) 0.305 (0.04)
ISIW-V Known 1.193 (0.11) 0.656 (0.13) 0.908 (0.04) 0.346 (0.06)
CvL 1.660 (0.12) 0.728 (0.16) 1.238 (0.06) 0.353 (0.06)
CvL.adaptive 1.639 (0.18) 0.714 (0.20) 1.098 (0.07) 0.362 (0.07)
INLA 1.686 (0.11) 0.783 (0.17) 1.204 (0.05) 0.355 (0.06)
diggle 1.616 (0.18) 0.712 (0.19) 1.112 (0.14) 0.355 (0.07)
ppl 1.664 (0.13) 0.730 (0.23) 1.168 (0.16) 0.361 (0.13)
scott 1.689 (0.11) 0.737 (0.20) 1.279 (0.06) 0.353 (0.06)
ISIW-PM Known 1.210 (0.10) 0.773 (0.18) 1.066 (0.06) 0.384 (0.08)
CvL 1.736 (0.12) 0.846 (0.21) 1.390 (0.07) 0.390 (0.09)
CvL.adaptive 1.614 (0.16) 0.766 (0.18) 1.175 (0.06) 0.395 (0.09)
INLA 1.764 (0.12) 1.006 (0.26) 1.331 (0.06) 0.392 (0.09)
diggle 1.593 (0.16) 0.774 (0.19) 1.136 (0.06) 0.382 (0.08)
ppl 1.723 (0.13) 0.804 (0.18) 1.238 (0.06) 0.380 (0.08)
scott 1.788 (0.12) 0.845 (0.19) 1.448 (0.07) 0.393 (0.09)
Table S2: Average RMSPE (standard deviation) for methods under the SCP simulation. Bolded values indicate the method with the lowest average RMSPE across all simulations for that setting.
N = 100 N = 800
Method Weights $\phi=0.02$ $\phi=0.15$ $\phi=0.02$ $\phi=0.15$
MLE 1.289 (0.06) 0.579 (0.10) 0.999 (0.04) 0.271 (0.03)
INLA-SLP 1.374 (0.31) 0.521 (0.08) 0.935 (0.03) 0.258 (0.02)
ISIW-V Known 1.228 (0.08) 0.551 (0.08) 0.889 (0.03) 0.270 (0.03)
CvL 1.280 (0.07) 0.566 (0.09) 0.975 (0.04) 0.268 (0.03)
CvL.adaptive 1.292 (0.09) 0.574 (0.09) 0.941 (0.05) 0.272 (0.03)
INLA 1.289 (0.07) 0.581 (0.10) 0.984 (0.04) 0.270 (0.03)
diggle 1.286 (0.08) 0.570 (0.10) 0.976 (0.04) 0.268 (0.03)
ppl 1.288 (0.06) 0.569 (0.09) 0.987 (0.04) 0.269 (0.03)
scott 1.285 (0.06) 0.568 (0.09) 0.995 (0.04) 0.269 (0.03)
ISIW-PM Known 1.232 (0.09) 0.596 (0.11) 0.900 (0.03) 0.281 (0.03)
CvL 1.287 (0.07) 0.612 (0.12) 1.017 (0.04) 0.281 (0.04)
CvL.adaptive 1.287 (0.10) 0.609 (0.12) 0.964 (0.04) 0.278 (0.03)
INLA 1.298 (0.07) 0.660 (0.14) 1.020 (0.04) 0.287 (0.04)
diggle 1.282 (0.08) 0.601 (0.11) 1.017 (0.04) 0.281 (0.04)
ppl 1.303 (0.06) 0.630 (0.12) 1.033 (0.04) 0.283 (0.04)
scott 1.296 (0.06) 0.628 (0.12) 1.043 (0.04) 0.285 (0.04)
Table S3: Average RMSPE (standard deviation) for methods under the Thomas process simulation. Bolded values indicate the method with the lowest average RMSPE across all simulations for that setting.
N = 100 N = 800
Method Weights $\phi=0.02$ $\phi=0.15$ $\phi=0.02$ $\phi=0.15$
MLE 1.684 (0.11) 0.822 (0.20) 1.324 (0.06) 0.376 (0.07)
INLA-SLP 1.214 (0.11) 0.749 (0.27) 0.983 (0.04) 0.324 (0.07)
ISIW-V Known 1.207 (0.11) 0.694 (0.15) 0.918 (0.06) 0.359 (0.06)
CvL 1.641 (0.14) 0.765 (0.18) 1.265 (0.06) 0.367 (0.07)
CvL.adaptive 1.627 (0.20) 0.752 (0.18) 1.121 (0.09) 0.375 (0.07)
INLA 1.629 (0.12) 0.812 (0.20) 1.216 (0.06) 0.369 (0.07)
diggle 1.580 (0.18) 0.766 (0.20) 1.137 (0.14) 0.367 (0.07)
ppl 1.682 (0.26) 0.909 (0.41) 1.278 (0.28) 0.412 (0.25)
scott 1.644 (0.13) 0.772 (0.24) 1.294 (0.06) 0.367 (0.07)
ISIW-PM Known 1.217 (0.08) 0.816 (0.21) 1.067 (0.06) 0.403 (0.09)
CvL 1.744 (0.15) 0.918 (0.28) 1.435 (0.08) 0.414 (0.11)
CvL.adaptive 1.603 (0.17) 0.816 (0.21) 1.182 (0.07) 0.409 (0.10)
INLA 1.720 (0.12) 1.040 (0.28) 1.337 (0.06) 0.411 (0.10)
diggle 1.558 (0.16) 0.830 (0.22) 1.168 (0.07) 0.397 (0.09)
ppl 1.617 (0.18) 0.816 (0.21) 1.239 (0.07) 0.397 (0.09)
scott 1.762 (0.13) 0.886 (0.22) 1.482 (0.08) 0.413 (0.10)
Table S4: Mean (SD) and Median (IQR) runtime in seconds over all simulations for methods under different sample sizes.
N = 100 N = 800
Method Variant Mean (SD) Median (IQR) Mean (SD) Median (IQR)
MLE 0.80 (0.26) 0.76 (0.34) 43.8 (13.8) 41.3 (17.2)
INLA-SLP 124.0 (71.8) 100.0 (87.3) 99.8 (39.3) 95.2 (37.3)
ILIW Known 4.80 (1.81) 4.47 (2.20) 45.8 (16.5) 43.3 (20.9)
CvL 4.14 (1.48) 3.85 (1.89) 34.4 (10.7) 32.7 (13.6)
CvL.adaptive 4.97 (1.75) 4.63 (2.11) 41.6 (13.7) 39.4 (17.5)
INLA 15.9 (4.51) 15.0 (5.39) 118.0 (41.4) 107.0 (51.0)
diggle 4.83 (1.93) 4.44 (2.24) 42.5 (17.3) 39.1 (22.1)
ppl 4.88 (1.89) 4.49 (2.16) 39.3 (16.2) 36.0 (16.8)
scott 4.08 (1.44) 3.79 (1.83) 34.3 (11.0) 32.6 (13.8)
PMLE Known 0.26 (0.06) 0.25 (0.06) 16.5 (2.4) 16.2 (3.3)
CvL 0.25 (0.05) 0.24 (0.05) 14.9 (2.1) 14.8 (2.8)
CvL.adaptive 0.59 (0.14) 0.55 (0.15) 16.6 (2.1) 16.4 (2.9)
INLA 9.61 (3.90) 8.40 (3.41) 97.4 (36.4) 86.3 (41.8)
diggle 0.28 (0.06) 0.27 (0.06) 16.4 (2.3) 16.2 (3.2)
ppl 0.65 (0.13) 0.64 (0.18) 15.7 (2.1) 15.6 (2.8)
scott 0.25 (0.05) 0.25 (0.05) 15.0 (2.3) 14.8 (3.3)
Table S5: Parameter estimates for the Galicia data.
Year Method $\mu$ $\sigma^{2}$ $\phi$ (km) $\tau^{2}$
1997 MLE 5.15 2.85 15.35 2.24
INLA-SLP 10.79 18.19 17.54 0.00007
ISIW-V 6.07 2.18 8.45 1.04
ISIW-PM 5.95 1.24 3.99 2.08
2000 MLE 2.24 1.40 8.90 0.000
INLA-SLP 2.27 1.36 10.2 0.00013
ISIW-V 2.24 1.40 8.91 0.000
ISIW-PM 2.15 1.37 9.47 0.016

### v2

@@
 ISIW is a two-stage approach wherein we estimate the sampling intensity of the observation locations then define intensity-based weights within a weighted likelihood adjustment.
-Prediction follows by substituting the adjusted parameter estimates within a kriging framework.
-A primary contribution was to implement ISIW by means of the Vecchia approximation, which provides large computational gains and improvements in predictive accuracy.
-Interestingly, we found that accurate parameter estimation had little correlation with predictive performance, raising questions about the conditions and parameter choices driving optimal implementation of kriging-based predictors under PS.
+Prediction follows by substituting the adjusted parameter estimates within kriging.
+We introduce an implementation of ISIW based on the Vecchia approximation, enabling computational gains while maintaining strong predictive accuracy.
+Interestingly, we found that ISIW outpredicts standard PS methods under misspecification of the sampling design, and that accurate parameter estimation had little correlation with predictive performance, raising questions about the conditions driving optimal implementation of kriging-based predictors under PS.
 Our work highlights the potential of ISIW to adjust for PS in an intuitive, fast, and effective manner.
-Keywords: model-based geostatistics, preferential sampling, kriging, point process, inverse weighting
+We illustrate these ideas on spatial prediction of lead concentrations measured through moss biomonitoring data in Galicia, Spain, and $\text{PM}_{2.5}$ concentrations from the U.S.
+EPA Air Quality System network in California.
+Keywords: inverse sampling intensity weighting, ISIW, model-based geostatistics, preferential sampling, kriging, point process
 1 Introduction
 The field of geostatistics comprises a set of methods to infer properties and predict unknown values (or functions) of a spatially continuous process $\{S(\textbf{x}):\textbf{x}\in\mathcal{D}\subseteq\mathbb{R}^{2}\}$ from a discrete set of observation locations, denoted $\textbf{X}=\{\textbf{x}_{1},\textbf{x}_{2},\ldots,\textbf{x}_{n}\}$.
-Model-based approaches estimated via maximum likelihood estimation (MLE) typically assume $S$ is a realization from a random stochastic process, and subsequent inference and prediction treat X as fixed (Diggle et al.,, 1998).
-This assumption further implies that the sampled locations are independent of the spatial process of interest, or in symbols, $[\textbf{X},S]=[\textbf{X}][S]$, where square brackets indicate the probability distribution.
+Model-based approaches estimated via maximum likelihood estimation (MLE) typically assume $S$ is a realization from a random stochastic process, and subsequent inference and prediction treat X as fixed (diggle_model-based_1998).
+This assumption further implies that the sampled locations are independent of the spatial process of interest, or in symbols, $\textbf{X},S]=[\textbf{X}][S]$, where square brackets indicate the probability distribution.
 Violations of this independence assumption are referred to as preferential sampling (PS).
-There is strong evidence to suggest neglecting PS can negatively impact both geostatistical inference and prediction (Diggle et al.,, 2010; Pati et al.,, 2011; Gelfand et al.,, 2012).
+Several studies have shown that neglecting PS can introduce substantial bias in geostatistical inference and prediction (diggle_geostatistical_2010; pati_bayesian_2011; gelfand_effect_2012).
 Early works in PS examined X under the lens of a marked point process.
-Schlather et al., (2004) designed two Monte Carlo tests to detect dependence between marks and their locations.
-Ho and Stoyan, (2008) introduced the classic PS adjustment model, referred to here as the shared latent process (SLP) model, and derived several of its properties.
-Full estimation and prediction procedures for the SLP model were formally established in the seminal work by Diggle et al., (2010).
+schlather_detecting_2004 designed two Monte Carlo tests to detect dependence between marks and their locations.
+ho_modelling_2008 introduced the classic PS adjustment model, referred to here as the shared latent process (SLP) model, and derived several of its properties.
+Full estimation and prediction procedures for the SLP model were formally established in the seminal work by diggle_geostatistical_2010.
 The SLP framework induces dependence between the spatial process of interest and the sampling locations through a shared variable approach.
@@
 Then the SLP model is as follows:
-$\begin{split}[\textbf{Y}|\textbf{X},S]\sim N_{n}(\mu\textbf{1}_{n}+S(\textbf{X}),\tau^{2}I_{n}),\\ [\textbf{X}|S]\sim IPP(\lambda),\\ \lambda(\textbf{x})=\exp(\alpha+\beta S(\textbf{x})),\\ S\sim GP(0,C_{\theta}(\cdot,\cdot)).\\ \end{split}$ (1)
-While the SLP framework has been extended to handle covariates in the mean model, non-Gaussian likelihoods, and multiple shared latent processes (Pati et al.,, 2011; Watson et al.,, 2019), for the purposes of illustration, we focus on the original formulation of a constant mean, Gaussian likelihood, and one shared latent process.
+[…]
+While the SLP framework has been extended to handle covariates in the mean model, non-Gaussian likelihoods, and multiple shared latent processes (pati_bayesian_2011; watson_general_2019), for the purposes of illustration, we focus on the original formulation of a constant mean, Gaussian likelihood, and one shared latent process.
 Under the SLP model, sampling locations are no longer assumed to be fixed and are instead a realization from a log Gaussian Cox process (LGCP).
-The latent process $S$ is assumed to follow a zero mean second order stationary Gaussian process (GP) and for the remainder of this paper, we further assume $C_{\theta}$ to be an isotropic Matérn covariance function.
+The latent process $S$ is assumed to follow a zero mean second order stationary Gaussian process (GP).
 The SLP takes its name from the presence of $S$ in both the mean model of Y and the intensity function of X.
@@
 Subsequent research on PS methods has remained largely faithful to the SLP model.
-Pati et al., (2011) extended the SLP model structure to a flexible Bayesian framework estimated by MCMC, and proved the posterior consistency of each of the mean, covariance and PS parameters under increasing domain asymptotics.
-Gelfand et al., (2012) also used Bayesian estimation, and introduced a framework to compare prediction surfaces under PS between different methods.
+pati_bayesian_2011 extended the SLP model structure to a flexible Bayesian framework estimated by MCMC, and proved the posterior consistency of each of the mean, covariance and PS parameters under increasing domain asymptotics.
+gelfand_effect_2012 also used Bayesian estimation, and introduced a framework to compare prediction surfaces under PS between different methods.
 A surprising finding from their simulation analysis was that inclusion of an informative covariate was not sufficient to correct for predictive bias.
 This empirical finding highlights the importance of the SLP model even when informative covariates explaining the dependence between X and $S$ are available.
-Dinsdale and Salibian-Barrera, (2019) significantly improved the computational efficiency of the SLP model relative to the previous simulation-based methods by means of the stochastic partial differential equations (SPDE) approach (Lindgren et al.,, 2011) combined with a Laplace approximation implemented in the Template Model Builder (TMB) R package (Kristensen et al.,, 2016).
-Watson et al., (2019) also fitted the SLP model using the SPDE approach but with an integrated nested Laplace approximation implemented in the R-INLA software (Rue et al.,, 2009), which enjoys comparable computational gains to the approach in Dinsdale and Salibian-Barrera, (2019).
+dinsdale_methods_2019 significantly improved the computational efficiency of the SLP model relative to the previous simulation-based methods by means of the stochastic partial differential equations (SPDE) approach (lindgren_explicit_2011) combined with a Laplace approximation implemented in the Template Model Builder (TMB) R package (kristensen_tmb_2016).
+watson_general_2019 also fitted the SLP model using the SPDE approach but with an integrated nested Laplace approximation implemented in the R-INLA software (rue_approximate_2009), which enjoys comparable computational gains to the approach in dinsdale_methods_2019.
 The authors further defined a framework to model PS spatio-temporal data and better emulate the evolution of X over time compared to the original SLP model defined in (1) above.
-Several other contributions extend the SLP framework to explore a rich set of important applications, including air pollution monitoring (Lee et al.,, 2011, 2015), species distribution modeling (Manceur and Kühn,, 2014; Fithian et al.,, 2015; Conn et al.,, 2017; Pennino et al.,, 2019; Gelfand and Shirota,, 2019; Fandos et al.,, 2021), disease surveillance (Rinaldi et al.,, 2015; Cecconi et al.,, 2016; Conroy et al.,, 2023), phylodynamic inference (Karcher et al.,, 2016), hedonic modelling (Paci et al.,, 2020), bivariate spatial data (Shirota and Gelfand,, 2022), spatially-varying PS (Amaral et al.,, 2023), optimal design under PS (Ferreira and Gamerman,, 2015), space-filling designs (Ferreira,, 2020; Gray and Evangelou,, 2023), a hypothesis test to detect PS in spatio-temporal data (Watson,, 2021), and exact Bayesian inference for the SLP model (Moreira and Gamerman,, 2022; Moreira et al.,, 2023).
-The improved predictive ability and parameter estimation within the SLP framework contribute to the popularity of model-based approaches accounting for dependence between spatial processes and their observation locations.
-While the benefits of the SLP structure and the ubiquity of PS data support its widespread adoption for spatial data analysis, the SLP model still contains significant limitations.
+Several other contributions extend the SLP framework to explore a rich set of important applications, including air pollution monitoring (lee_constructing_2011; lee_impact_2015), species distribution modeling (manceur_inferring_2014; fithian_bias_2015; conn_confronting_2017; pennino_accounting_2019; gelfand_preferential_2019; fandos_dynamic_2021), disease surveillance (rinaldi_sheep_2015; cecconi_preferential_2016; conroy_shared_2023), phylodynamic inference (karcher_quantifying_2016), hedonic modelling (paci_spatial_2020), bivariate spatial data (shirota_preferential_2022), spatially-varying PS (amaral_model-based_2023), optimal design under PS (ferreira_optimal_2015), space-filling designs (ferreira_geostatistics_2020; gray_design_2023), a hypothesis test to detect PS in spatio-temporal data (watson_perceptron_2021), and exact Bayesian inference for the SLP model (moreira_analysis_2022; moreira_presence-only_2024).
+The SLP’s flexibility in addressing sampling bias has driven the popularity of model-based approaches that account for the inherent dependence between spatial processes and their observation locations.
+Despite these advantages, the SLP framework contains significant limitations that complicate its application.
 Few software packages and out-of-the-box implementations exist for SLP estimation, often coming in the form of custom INLA or MCMC sampler code.
-Furthermore, the SLP model is challenging to integrate with modern spatial approximation techniques, many of which have become the standard in spatial data analysis due to their scalability and feasibility, offering computationally efficient alternatives to the MLE while maintaining high accuracy (Heaton et al.,, 2019).
+Furthermore, the SLP model is challenging to integrate with modern spatial Gaussian process approximation techniques, many of which have become the standard in spatial data analysis due to their scalability and feasibility, offering computationally efficient alternatives to the MLE while maintaining high accuracy (heaton_case_2019).
 Among these methods, only the SPDE approach has a well-documented implementation of a PS solution, and incorporating the SLP model into any other method would take a substantial effort.
-The SLP model requires joint evaluation of both the response and observation likelihood, not a trivial task.
-Prediction based on the SLP model also can suffer from high computational cost.
-While parameter estimation for the SLP model may scale well with the SPDE approach, prediction depends on summarizing the posterior distribution of $[S|\textbf{X},\textbf{Y}]$ for each prediction point, which can be infeasible even on a moderately sized grid for both the Laplace approximation and MCMC.
+The SLP model requires the nontrivial task of jointly modeling both the response and observation likelihood.
+Prediction based on the SLP model can also suffer from high computational cost.
+While parameter estimation for the SLP model may scale well with the SPDE approach, prediction depends on summarizing the posterior distribution of $S|\textbf{X},\textbf{Y}]$ for each prediction point, which can be infeasible even on a moderately sized grid for both the Laplace approximation and MCMC.
 An alternative strategy to the SLP is to incorporate dependence through the sampling intensities $\lambda$ of the observations, rather than the entire likelihood of X.
-These methods have been developed to much less fanfare compared to the SLP, and it remains unclear how well they work in geostatistical applications.
-The two main uses of sampling intensities are as a covariate in the mean model of Y or as inverse weights in a likelihood adjustment (Vedensky et al.,, 2023).
-A potential advantage of sampling intensity methods is their robustness to the form of PS relative to the SLP framework, which addresses a very specific type of clustering process.
-The biggest impediment to the use of the sampling intensity is the need to estimate this intensity from the observed locations, the driving reason such approaches have been sidelined in favor of full model-based approaches.
+The two main potential uses of sampling intensities are as a covariate in the mean model of Y or as inverse weights in a likelihood adjustment (vedensky_look_2023).
+It remains unclear how well these methods work in geostatistical applications.
+A potential advantage of sampling intensity methods is their robustness to the form of PS relative to the SLP framework, which requires full parametric specification of the likelihood of $\mathbf{X}$.
+The biggest impediment to the use of the sampling intensity is the need to estimate this intensity from the observed locations.
 Unlike in survey methodology where survey weights are fixed and known, sampling weights in geostatistical applications are largely unknown to the investigator and must be estimated.
-Unfortunately, nonparametric kernel smoothing estimators of the intensity (Diggle,, 1985; Berman and Diggle,, 1989) have few theoretical guarantees and are not consistent for the true intensity (Guan,, 2008).
-Even so, there is preliminary evidence that methods using nonparametrically estimated sampling intensities can still mitigate the effects of PS and, in some scenarios, can predict a surface better than the SLP approach in practice.
-Reich and Fuentes, (2012) noted how the Bayesian SLP model parameters could be estimated without MCMC by replacing the shared latent process term in the mean model with some function of the sampling intensity $g(\lambda)$.
-Since the dependence between X and $S$ is completely contained within $g(\lambda)$, the locations X can once again be treated as fixed, so long as $\lambda$ is estimated well.
-They show improved prediction using the sampling intensity covariate in a 1-dimensional kriging example.
-In another example, Zidek et al., (2014) provide an approach using estimated weights for air pollution monitoring.
+Unfortunately, nonparametric kernel smoothing estimators of the intensity (diggle_kernel_1985; berman_estimating_1989) have few theoretical guarantees and are not consistent for the true intensity without knowledge of spatial covariates (guan_consistent_2008).
+Even so, there is preliminary evidence that methods using nonparametrically estimated sampling intensities can still mitigate the effects of PS.
+mateu_accounting_2012 noted how the Bayesian SLP model parameters could be estimated without MCMC by replacing the shared latent process term in the mean model with some known function of the sampling intensity $g(\lambda)$.
+In another example, zidek_reducing_2014 provide an approach using estimated weights for air pollution monitoring.
 Instead of working in a continuous study region, however, these authors considered a finite superpopulation of possible sampling sites and the probability of selecting any site over time was modeled by a logistic regression.
 The estimated inverse probabilities were then used as weights in a Horvitz-Thompson style design-based estimator for unbiased estimation of parameters under PS.
-In a third example, Schliep et al., (2023) also considered a finite superpopulation and similarly deviated from traditional model-based geostatistics by using estimated sampling intensity weights to recover the estimated parameter values had the model been fit on the superpopulation, rather than estimate the parameters of the spatial process $S$.
-These authors also derived a kriging variance estimate adjusted for PS.
-Finally, Vedensky et al., (2023) conducted a simulation experiment comparing a univariate marginal composite likelihood (CL) weighted by inverse sampling intensities estimated by the MASS::kde2d function in R to the SLP and unweighted CL, and showed improved performance compared to no adjustment.
+schliep_correcting_2023 also considered a finite superpopulation and similarly deviated from traditional model-based geostatistics by using estimated sampling intensity weights to recover the estimated parameter values and kriging variance had the model been fit on the superpopulation, rather than estimate the parameters of the spatial process $S$.
+Finally, vedensky_look_2023 conducted a simulation experiment comparing a univariate marginal composite likelihood (CL) weighted by inverse sampling intensities estimated by the MASS::kde2d function in R to the SLP and unweighted CL, and showed improved performance compared to no adjustment.
 We will refer to methods using inverse sampling intensities as weights in a weighted likelihood adjustment as inverse sampling intensity weighting (ISIW).
-Despite preliminary investigations into ISIW for PS adjustment, specifics regarding the effectiveness of such approaches within model-based geostatistics remains largely unknown.
+Despite preliminary investigations into ISIW for PS adjustment, specifics regarding the effectiveness of such approaches within model-based geostatistics remain largely unknown.
 In particular, it is unclear whether they reliably estimate both mean and covariance parameters for a latent spatial process on a 2D continuous surface.
@@
 Finally, the robustness of ISIW and the SLP to misspecification has not been well-studied.
-In the sections below, we provide an expanded evaluation of ISIW methods, comparing the performance of MLE, SLP, and ISIW applied to the Vecchia and pairwise marginal CLs across multiple random fields and PS designs.
+In the sections below, we provide an expanded evaluation of ISIW methods, comparing the performance of MLE, SLP, and ISIW across multiple random fields and PS designs.
 Our work proceeds as follows.
@@
 In Section 4, we conduct a comparison analysis of the MLE, SLP, and ISIW through a set of simulation studies.
-In Section 5 we apply the same methods to the famous Galicia moss dataset which has been the dataset of choice when investigating PS.
+In Section 5 we apply the same methods to the famous Galicia moss dataset which has been the dataset of choice when investigating PS, as well as a dataset of air pollution measurements taken from the Air Quality System (AQS) in California.
 In Section 6 we finish with final remarks and discuss the implications of our work, and outline future directions for continuing investigation.
 2 Model-based geostatistics
-2.1 Estimation
-2.1.1 Maximum likelihood estimation
+2.1 Maximum likelihood estimation
 Under non-preferential sampling (NPS), we define a model for the Gaussian observations Y and underlying Gaussian process $S$ following the SLP framework in (1), but we drop the point process likelihood for X.
 The standard geostatistical model under NPS follows
-$\begin{split}[\textbf{Y}|\textbf{X},S]\sim N_{n}(\mu\textbf{1}_{n}+S(\textbf{X}),\tau^{2}I_{n}),\\ S\sim GP(0,C_{\boldsymbol{\theta}}(\cdot,\cdot)).\\ \end{split}$
+[…]
 We further assume the covariance function $C_{\theta}$ follows a stationary isotropic Matérn covariance defined as
@@
 Because any finite collection of random variables corresponding to point observations of $S$ follows a multivariate Gaussian distribution and X is treated as fixed, the observed data likelihood is then defined by
-$\begin{split}[\textbf{Y},\textbf{X}]\propto N_{n}(\mu\textbf{1}_{n},\Sigma_{n}(\theta)+\tau^{2}I_{n}),\\ \end{split}$ (2)
+[…]
 where $\Sigma_{n}$ is the covariance matrix with $ij$th entry equal to $C_{\theta}(\lVert\textbf{x}_{i}-\textbf{x}_{j}\rVert)$.
 Estimation proceeds by optimization of the likelihood with respect to the parameter vector $\boldsymbol{\psi}$.
-Justification for geostatistical inference via the MLE derives from spatial large sample theory.
-Two main frameworks have dominated asymptotics for geostatistical estimators: increasing domain asymptotics, and fixed domain (or infill) asymptotics.
-Increasing domain asymptotics assume that as $n\to\infty$, the study region $\mathcal{D}$ expands, ensuring a minimum separation distance between observation locations and a fixed density of observations per unit area.
-Fixed domain asymptotics, on the other hand, keep $\mathcal{D}$ fixed while increasing $n$, leading to a growing observation density and vanishing minimum distance between observations (i.e., observations occur closer together as the sample size increases within the fixed study area).
-Under increasing domain asymptotics, the MLE for $\boldsymbol{\psi}$ is consistent and asymptotically normal (AN) (Mardia and Marshall,, 1984; Bachoc,, 2014, 2020).
-Results are more challenging under fixed domain asymptotics due to the inclusion of increasingly close observations of a spatially correlated process.
-Specifically, for a Matérn Gaussian process with known $\nu$, only a subcomponent of the covariance parameters known as the microergodic parameter is consistently estimable and asymptotically normal (Zhang,, 2004; Kaufman and Shaby,, 2013).
-We refer to this parameter as $\kappa$, given by $\sigma^{2}/\phi^{2\nu}$ for Matérn covariance functions.
-The nugget $\tau^{2}$ is also estimable (Tang et al.,, 2021) while the variance $\sigma^{2}$ and range $\phi$ are not (Zhang,, 2004).
-Although the smoothness parameter $\nu$ is theoretically estimable, it is numerically challenging to estimate and is conventionally treated as fixed (Loh et al.,, 2021).
-Fixed effect coefficients in the mean are typically non-estimable (Wang et al.,, 2020), except under specific smoothness conditions on the covariates (Yu,, 2022; Bolin and Wallin,, 2024; Gilbert et al.,, 2024).
 The main drawback for practical use of the MLE is its computational burden.
@@
 Recent advancement in modern spatial statistics has focused on GP parameter estimation using approximations which massively reduce computational burden in exchange for marginal decreases in efficiency.
-We explore composite likelihood and the Vecchia approximation, two approaches which have been widely adopted in geostatistics and are particularly conducive to weighting.
-2.1.2 Composite likelihood
-Sticking with our previous notation, let $\textbf{Y}=(Y_{1},Y_{2},...,Y_{n})^{\top}\in\mathbb{R}^{n}$ be a $n\times 1$ random vector with probability density $f(\textbf{y};\boldsymbol{\psi})$ for unknown $r$-dimensional parameter vector $\boldsymbol{\psi}\in\Psi\subseteq\mathbb{R}^{r}$.
-Define $\{\mathcal{A}_{1},...,\mathcal{A}_{K}\}$ to be a set of $K$ marginal or conditional events with associated likelihood $\mathcal{L}_{k}(\theta;\textbf{y})\propto f(\textbf{y}\in\mathcal{A}_{k};\boldsymbol{\psi})$.
-The log composite likelihood (CL) is the weighted sum of the event-specific log likelihoods
-$\log\mathcal{L}_{C}(\boldsymbol{\psi};\textbf{y}):=\sum_{k=1}^{K}w_{k}\log\mathcal{L}_{k}(\boldsymbol{\psi};\textbf{y}),$
-where $\textbf{w}:=(w_{1},...,w_{K})$ is a vector of weights, not necessarily non-negative and the density $f$ follows that defined in (2).
-The maximum CL estimate is defined as $\hat{\boldsymbol{\psi}}_{C}:=\text{argmax}_{\boldsymbol{\psi}}\log\mathcal{L}_{C}(\boldsymbol{\psi};\textbf{y})$.
-One common choice of CL is the univariate marginal where the log likelihood is the sum of the log marginal density of each observation.
-$\log\mathcal{L}_{UM}(\boldsymbol{\psi};\textbf{y})=\sum_{i=1}^{n}w_{i}\log f(y_{i};\boldsymbol{\psi}).$
-Typically, the use of this univariate marginal in geostatistics is limited because it ignores dependencies between observations and thus cannot estimate covariance parameters.
-For this reason, pairwise CLs have been much more popular (Varin et al.,, 2011).
-Bevilacqua and Gaetan, (2015) compared the efficiency of three different pairwise CL estimators: pairwise marginal (PMLE), pairwise conditional (PCMLE), and pairwise difference (PDMLE) CLs.
-These authors concluded that the PMLE outperformed all other pairwise CLs and recommended its use over the PCMLE and PDMLE.
-The likelihoods for each pairwise CL are
-$\begin{split}\log\mathcal{L}_{PM}(\boldsymbol{\psi};\textbf{y})=\sum_{i<j}w_{ij}\log f(y_{i},y_{j};\boldsymbol{\psi}),\\ \log\mathcal{L}_{PC}(\boldsymbol{\psi};\textbf{y})=\sum_{i\neq j}w_{ij}\log f(y_{i}|y_{j};\boldsymbol{\psi}),\\ \log\mathcal{L}_{PD}(\boldsymbol{\psi};\textbf{y})=\sum_{i<j}w_{ij}\log f(y_{i}-y_{j};\boldsymbol{\psi}).\end{split}$ (3)
-These authors also proved, under increasing domain asymptotics, the consistency and asymptotic normality of $\boldsymbol{\hat{\psi}}_{C}$ estimated from (3).
-Bachoc et al., (2019) proved the same properties for the PMLE and PCMLE in the one-dimensional setting under fixed domain asymptotics.
-However, the true utility of CL lies in its computational speed and robustness to misspecification.
-While the MLE requires a $O(n^{3})$ matrix inversion and specification of the joint density, pairwise CL consists only of $O(n^{2})$ terms and only requires correct specification of second order densities.
-The computational efficiency of pairwise CL can be further enhanced by using the weights w to only include pairwise observations within a certain distance $d$ apart (Bevilacqua and Gaetan,, 2015).
-In later sections, we will use the PMLE as a basis for ISIW due to its superior performance over other pairwise likelihoods.
-In particular, we do not consider the univariate marginal or PDMLE because they do not directly estimate the entirety of $\boldsymbol{\psi}$.
-In the next section, we discuss the Vecchia approximation, an alternative likelihood approach that achieves even greater computational and statistical efficiency than the CL.
-2.1.3 Vecchia approximation
-The Vecchia approximation is a specific case of CL based on the observation that the joint distribution can be decomposed as the product of conditional distributions.
+In the next section, we discuss the Vecchia approximation, one such alternative likelihood approach that achieves high computational and statistical efficiency.
+2.2 Vecchia approximation
+The Vecchia approximation is a specific case of composite likelihood (varin_overview_2011) based on the observation that the joint distribution of a random vector can be decomposed as the product of conditional distributions.
 Let $\textbf{Y}=(Y_{1},...,Y_{n})^{\top}\in\mathbb{R}^{n}$ and $p:\{1,...,n\}\to\{1,...,n\}$ be a permutation mapping which reorders Y.
+Here $f(\cdot;\boldsymbol{\psi})$ denotes the probability density of the corresponding subvector of $\mathbf{Y}$ under parameter vector $\boldsymbol{\psi}$.
 We define the history of variable $Y_{j}$ as a random subvector $\textbf{Y}_{h(j)}$ where $h(j)=\{l\in\{1,...,n\}:p(l)<p(j)\}$.
-The joint density of Y can then be refactored as
+The joint density of Y evaluated at a fixed value $\mathbf{y}$ can then be refactored as
 $f(\textbf{y};\boldsymbol{\psi})=f(y_{p(1)};\boldsymbol{\psi})\prod_{i=2}^{n}f(y_{p(i)}|\textbf{y}_{h(i)};\boldsymbol{\psi}).$
@@
 The Vecchia approximation can also be written in a weighted CL form,
-$\begin{split}\log\mathcal{L}_{V}(\boldsymbol{\psi};\textbf{y})&=w_{1}\log f(y_{p(1)};\boldsymbol{\psi})\\ &+\sum_{i=2}^{n}w_{1i}f(y_{p(i)},\textbf{y}_{q(i)};\boldsymbol{\psi})\\ &-\sum_{i=2}^{n}w_{2i}\log f(\textbf{y}_{q(i)};\boldsymbol{\psi}).\end{split}$ (5)
+$\begin{split}\log\mathcal{L}_{V}(\boldsymbol{\psi};\textbf{y})&=w_{1}\log f(y_{p(1)};\boldsymbol{\psi})\\ &+\sum_{i=2}^{n}w_{1i}\log f(y_{p(i)},\textbf{y}_{q(i)};\boldsymbol{\psi})\\ &-\sum_{i=2}^{n}w_{2i}\log f(\textbf{y}_{q(i)};\boldsymbol{\psi}).\end{split}$ (3)
 While simple in concept, the Vecchia approximation requires careful selection of three key hyperparameters: 1) the size of the conditioning set, 2) which variables to include in each conditioning set, and 3) the ordering of the variables as determined by $p$.
 We denote the maximum size of any $q(i)$ as $m$.
-In all later sections, we choose $m=20$ based on a clear case of diminishing returns in inference and prediction for $m>20$ observed empirically in Datta et al., (2016).
-Default settings in two implementations of the Vecchia approximation (GPvecchia and GpGp R packages) are $m=20$ and $m=30$, respectively (Katzfuss and Guinness,, 2021; Guinness,, 2021).
-For the choice of which variables to include in the neighborhood $q(i)$, we follow the recommendation given in Vecchia, (1988) by choosing the nearest neighbors to $y_{p(i)}$ measured by Euclidean distance.
-Ordering for the Vecchia approximation requires additional specification in multidimensional settings due to the lack of natural ordering in observations of a spatial point process.
-We use the maxmin ordering scheme, which picks a location $p(i)$ sequentially by maximizing the distance to the nearest point in $\{y_{p(1)},y_{p(2)},...,y_{p(i-1)}\}$ and has been shown to approximate the true distribution better than other ordering methods (Guinness,, 2018; Katzfuss and Guinness,, 2021).
-Theoretical justification of the Vecchia approximation in spatial statistics is relatively light compared to that of the MLE and CL.
-Typical approaches to proving fixed domain asymptotics fail due to the loss of stationarity in the parent process introduced by the Vecchia approximation (Zhang et al.,, 2024).
-However, it is impossible to deny its excellent empirical performance in applications (Heaton et al.,, 2019).
-A key advantage the Vecchia approximation possesses over typical CL is that (4) and (5) correspond to a valid joint probability distribution for Y.
-As a result, the Kullback-Leibler (KL) divergence of the Vecchia approximation with respect to the true distribution can be computed and has been shown to be a nonincreasing function of $m$ when the conditioning sets $q(i)$ are chosen as the nearest neighbors (Katzfuss et al.,, 2020; Huser et al.,, 2023).
-The Vecchia approximation achieves a good balance of computational complexity and efficiency.
-It clearly outperforms other CL methods in terms of what order correlations are being accounted for, scalability, and probabilistic validity.
-This is the main reason for considering its use in ISIW methods compared to CL, a point we expand on in later sections.
-2.2 Prediction
+In all later sections, we choose $m=20$ based on a clear case of diminishing returns in inference and prediction for $m>20$ observed empirically in datta_hierarchical_2016.
+Default settings in two implementations of the Vecchia approximation (GPvecchia and GpGp R packages) are $m=20$ and $m=30$, respectively (katzfuss_general_2021; guinness_gaussian_2021).
+For the choice of which variables to include in the neighborhood $q(i)$, we follow the recommendation given in vecchia_estimation_1988 by choosing the nearest neighbors to $y_{p(i)}$ measured by Euclidean distance.
+Ordering for the Vecchia approximation requires additional specification in 2D settings due to the lack of natural ordering in observations of a multidimensional point process.
+We use the maxmin ordering scheme, which picks a location $p(i)$ sequentially by maximizing the distance to the nearest point in $\{y_{p(1)},y_{p(2)},...,y_{p(i-1)}\}$ and has been shown to approximate the true distribution better than other ordering methods (guinness_permutation_2018; katzfuss_general_2021).
+The Vecchia approximation strikes a favorable balance between computational efficiency and statistical accuracy.
+Compared with other composite likelihood approaches, it accommodates higher-order dependence, scales well to large datasets, and corresponds to a valid joint density.
+This is the main reason for considering its use in ISIW methods compared to univariate and pairwise composite likelihoods from previous studies.
+2.3 Kriging
 For many spatial analyses, the main goal is to predict values at unobserved locations.
-We focus on spatial prediction via kriging, or Gaussian process regression, a function approximation technique with point and variance estimate defined as
+A standard spatial prediction tool is kriging, or Gaussian process regression, a function approximation technique with point and variance estimate defined as
 $\hat{S}(\textbf{X}_{0})$ $=\mu+C_{\theta}(\textbf{X}_{0},\textbf{X}_{n})^{\top}\boldsymbol{\Sigma}_{n}(\theta)^{-1}(\textbf{Y}-\mu\textbf{1}_{n}),$ (4)
-$\operatorname{Var}(\hat{S}(\textbf{X}_{0}))$ $=C_{\theta}(\textbf{X}_{0},\textbf{X}_{0})-C_{\theta}(\textbf{X}_{0},\textbf{X}_{n})^{\top}\boldsymbol{\Sigma}_{n}^{-1}(\theta)C_{\theta}(\textbf{X}_{0},\textbf{X}_{n}),$
+$\operatorname{Var}\big\{\hat{S}(\textbf{X}_{0})\big\}$ $=C_{\theta}(\textbf{X}_{0},\textbf{X}_{0})-C_{\theta}(\textbf{X}_{0},\textbf{X}_{n})^{\top}\boldsymbol{\Sigma}_{n}^{-1}(\theta)C_{\theta}(\textbf{X}_{0},\textbf{X}_{n}),$
 where $\textbf{X}_{0}$ represents unobserved locations to be predicted.
-In practice, since the parameters are unknown, all parameters in (6) are replaced with estimates obtained from any of the aforementioned or other estimation methods.
-Kriging has several properties that make it an effective predictor of spatial surfaces.
-As the best linear unbiased predictor (BLUP), it minimizes mean squared prediction error among all unbiased predictors (Cressie and Wikle,, 2015).
-However, it is important to note that this optimality no longer holds under preferential sampling, where even the true values of $\boldsymbol{\psi}$ used in the kriging equations yield suboptimal predictions (Dinsdale and Salibian-Barrera,, 2019).
-Stein established the asymptotic efficiency of kriging under various robustness conditions (Stein,, 1988; Stein, 1990a, ; Stein, 1990b, ; Stein,, 1993), while Wang et al., (2020) showed that kriging’s prediction error vanishes under a uniform metric.
-Additionally, Putter and Young, (2001) demonstrated that the difference between predictions using kriging with estimated and the true parameters is asymptotically negligible if the joint Gaussian distributions of the spatial process under the true and estimated covariance functions are contiguous almost surely.
-Given these results, one could argue that the geostatistical model in (2) is more appropriate for spatial interpolation rather than inference.
-3 Inverse sampling intensity weighting
-We next describe estimation and prediction procedures of ISIW for PS adjustment.
-ISIW is a two-stage approach, wherein we first estimate the sampling intensity at each of the observation locations X, and, in the second stage, we input the (estimated) inverse sampling intensities as weights into a weighted likelihood adjustment.
-The resulting adjusted parameter estimates are then substituted in the kriging equations for prediction.
-3.1 Estimation of sampling intensity
-The essence of ISIW is to account for the dependence between the response and observation locations by using the vector of estimated sampling intensities instead of the more complex full likelihood of the observation process.
-This approach assumes that the locations contain sufficient information about the spatial dependence structure of the observation locations to enable proper adjustment.
-Table 1: Point process intensity function estimators considered in the simulation study.
-Method Reference Bandwidth Selection
-diggle Diggle, (1985) Least-squares cross-validation
-scott Scott, (1992) Rule-of-thumb based on normal reference density
-ppl Loader, (1999) Likelihood cross-validation (leave-one-out)
-CvL Cronie and Van Lieshout, (2018) Maximum likelihood cross-validation
-CvL.adaptive van Lieshout, (2021) Adaptive bandwidth based on local point density
-R-INLA Simpson et al., (2016) LGCP model (non-kernel-based)
+In practice, since the parameters are unknown, all parameters in (4) are replaced with estimates obtained from any of the aforementioned estimation methods.
+It can be shown under the assumption of NPS that kriging is the best linear unbiased predictor (BLUP) (cressie_statistics_2015).
+However, this optimality no longer holds under preferential sampling, where even the true values of $\boldsymbol{\psi}$ used in the kriging equations yield suboptimal predictions (dinsdale_methods_2019).
+This is because the optimal predictor is the conditional expectation $\mathbb{E}\big\{S(\mathbf{X}_{0})\mid\mathbf{Y},\mathbf{X}\big\}$, whose distribution is no longer Gaussian as in the fixed case.
+Nevertheless, kriging can still be used to generate predictions under these non-standard conditions.
+One potential approach is to solve for the kriging weights under the model defined in (1).
+However this optimization does not admit a closed-form expression.
+We therefore turn to ISIW as a practical approach for finding GP mean and covariance parameters that yield an approximation to the optimal linear kriging predictor under PS, even though these parameters may not coincide with the true latent random field.
+3 Inverse sampling intensity weighting (ISIW)
+We now describe the ISIW procedure for PS adjustment.
+ISIW is a two-stage approach, wherein we first estimate the sampling intensity at each of the observation locations X.
+In the second stage, we input the (estimated) inverse sampling intensities as weights into a weighted likelihood adjustment for a chosen composite likelihood.
+For spatial interpolation, the resulting adjusted parameter estimates are then substituted in the kriging equations in (4).
+3.1 Sampling intensity estimation
+The essence of ISIW is to account for the dependence between the response and observation locations by using the vector of estimated sampling intensities instead of the full likelihood of the observation process.
+The heuristic motivation for ISIW is to reweight observations by inverse sampling intensity to approximate a representative, non-preferential spatial sample of the study region.
 Methods to estimate a spatially varying intensity of a point process can be divided into parametric and nonparametric approaches.
 Domain knowledge can inform the parametric form of the intensity either through choice of model or covariates, but oftentimes this information is unavailable and nonparametric estimation is required.
-The nonparametric approaches follow the kernel smoothing approach discussed in Diggle, (1985).
+The nonparametric approaches follow the kernel smoothing approach discussed in diggle_kernel_1985 which we will refer to as kernel intensity estimators (KIE).
 Let $k$ be a $d$-dimensional kernel function from $\mathbb{R}^{d}\to\mathbb{R}^{+}$, which is a symmetric probability density function.
-Given a bandwidth size $h>0$ and edge correction factor $w_{h}(\cdot,\cdot)$, the nonparametric kernel smoothing estimator of the intensity function is given by
-$\hat{\lambda}(\textbf{x};h)=h^{-d}\sum_{\textbf{y}\in\textbf{X}\cap\mathcal{D}}k\left(\frac{\lVert\textbf{x}-\textbf{s}\rVert}{h}\right)w_{h}(\textbf{x},\textbf{s})^{-1},\textbf{x}\in\mathcal{D}.$
-The key hyperparameter for nonparametric intensity estimation is the bandwidth size $h$.
+Given a bandwidth size $h>0$ and edge correction factor $w_{h}(\cdot,\cdot)$, the KIE estimate at a point $\mathbf{x}$ is given by
+$\hat{\lambda}(\textbf{x};h)=h^{-d}\sum_{\textbf{s}\in\textbf{X}\cap\mathcal{D}}k\left(\frac{\lVert\textbf{x}-\textbf{s}\rVert}{h}\right)w_{h}(\textbf{x},\textbf{s})^{-1},\textbf{x}\in\mathcal{D}.$
+The key hyperparameter for KIE is the bandwidth size $h$.
 Bandwidth selection is a well-studied problem, with methods ranging from high bandwidth, smooth intensity estimators to low bandwidth, flexible intensity estimators.
 Each has its place in the bias-variance tradeoff, with higher bandwidths exhibiting higher bias with lower variance and lower bandwidths vice versa.
-We experiment with several bandwidth selection strategies implemented in the spatstat R package (Table 1).
-The effectiveness of any ISIW approach for PS should improve the closer estimated weights $\lambda(\textbf{x})^{-1}$ are to the true inverse intensities.
-In theory, a parametric LGCP model should estimate the true intensity better than any nonparametric estimator if X follows a LGCP, especially with multiple realizations.
-However, it is known that it is impossible to distinguish between a single realization from an inhomogenous Poisson process with deterministic intensity $\lambda(\textbf{x})$ and one from a stationary Cox process whose conditional intensity coincides with $\lambda$ (Gelfand et al.,, 2010).
-We consider a LGCP method introduced in Simpson et al., (2016) estimated by R-INLA to verify whether a known parametric model shows improvement over a nonparametric estimator.
-Once the sampling intensities are estimated at every observation location, we calculate weights as the inverse of the sampling intensities and then normalize these to sum up to the number of observations $n$.
-The estimated weight for the $i$th observation becomes
-$\hat{w}_{i}=\hat{w}(\textbf{x}_{i})=n*\frac{\hat{\lambda}(\textbf{x}_{i})^{-1}}{\sum_{\textbf{s}\in\textbf{X}}\hat{\lambda}(\textbf{s})^{-1}}.$ (7)
-3.2 Defining the likelihood
+We experiment with several bandwidth selection strategies implemented in the spatstat R package (Table S1).
+While KIEs provide valid estimates of the first-order intensity based solely on a realization of the point pattern, a potential limitation for PS adjustment is that they do not incorporate information from the observed responses $\mathbf{Y}$.
+Under the SLP model in (1), for example, the sampling intensity depends on the latent process, which also drives the observed response values.
+Incorporating both response values and spatial locations may therefore yield improved estimates of the sampling intensity relative to approaches based only on the locations.
+To explore this idea, we study an extension of KIE that incorporates the response through a spatial covariate using the kernel-based approach of baddeley_nonparametric_2012.
+We also considered a random forest–based intensity estimation method proposed by biscio_nonparametric_2025; however, its results were very similar to those obtained with KIE so we do not present them here.
+Specifically, we use a predicted version of $\mathbf{Y}$ as an additional covariate in the intensity estimation.
+Because KIEs with spatial covariates require covariate values to be available at all grid locations, we obtain this covariate by kriging using parameters estimated via MLE.
+Although this predicted surface is biased, as has been extensively discussed, PS induces systematic discrepancies between the true process and the estimated surface.
+If such structure is present, the nonparametric relationship between the covariate and the sampling intensity may still be learnable.
+We refer to this approach as “ISIW KIE COV” and the approach using just the locations as “ISIW KIE.”
+As with other inverse weighting procedures, a key concern for ISIW is the presence of extreme weights.
+Observation locations in sparsely sampled regions, which occur more frequently under PS designs, can receive very large weights in the resulting likelihood and may lead to numerical instability during optimization.
+Two commonly used approaches from the inverse probability weighting literature are trimming and winsorization.
+Trimming removes observations deemed to be outliers, whereas winsorization preserves the sample size by capping extreme values at a specified threshold.
+We adopt winsorization to mitigate the impact of extreme weights, as trimming observations may discard too many informative locations, thereby weakening the effectiveness of the PS adjustment.
+Our winsorization procedure proceeds as follows.
+First, we estimate the sampling intensities and take their inverse.
+An upper quantile is then selected as the winsorization threshold, and all values exceeding this threshold are set equal to it.
+Finally, the weights are normalized to sum to the total number of observations $n$.
+Denote the $\tau$-level empirical quantile of the estimated inverse sampling intensities $\hat{\lambda}^{-1}$ as $W_{\tau}$.
+The resulting estimated weight for the $i$th observation is
+$\hat{w}_{i}=n\cdot\frac{\min\{\hat{\lambda}(\mathbf{x}_{i})^{-1},W_{\tau}\}}{\sum_{\mathbf{s}\in\mathbf{X}}\min\{\hat{\lambda}(\mathbf{s})^{-1},W_{\tau}\}}.$ (5)
+3.2 Weighted likelihood and estimation
 ISIW applies weights proportional to the inverse sampling intensity to each event in a likelihood factored as a product of densities.
-The only known examples of ISIW in the PS literature include a pairwise difference (Schliep et al.,, 2023) and univariate marginal CL (Vedensky et al.,, 2023).
-The pairwise difference CL cannot estimate $\mu$ directly, limiting ISIW’s impact on PS adjustment of the mean, while the univariate marginal CL cannot estimate covariance parameters $\theta$.
-To address these gaps, we propose two new ISIW estimators for simultaneous mean and covariance estimation: the ISIW pairwise marginal (ISIW-PM) and ISIW Vecchia (ISIW-V) estimator.
-Let the weights be defined as in (7).
-Then the ISIW-PM follows
-$\begin{split}\log\mathcal{L}_{WPM}(\boldsymbol{\psi};\textbf{y})=\sum_{i<j}w_{i}w_{j}\log f(y_{i},y_{j};\boldsymbol{\psi}).\end{split}$
-For the ISIW-V, we initially considered weighting the likelihood as follows
+While such weighting is straightforward in settings with independent observations, the appropriate weighting scheme is less clear in the presence of spatial dependence and correlation among observations.
+Existing applications of ISIW in the PS literature include univariate marginal (vedensky_look_2023) and pairwise difference CLs (schliep_correcting_2023).
+These approaches yield intuitive weighting schemes by enforcing independence across subsets of events in the joint density.
+For reference, the weighted univariate marginal CL follows
+$\begin{split}\log\mathcal{L}_{WM}(\mu;\textbf{y})=\sum_{i=1}^{n}w_{i}\log f(y_{i};\mu),\end{split}$
+while the pairwise difference CL follows
+$\begin{split}\log\mathcal{L}_{WD}(\theta;\textbf{y})=\sum_{i=1}^{n}\sum_{j=i+1}^{n}w_{i}w_{j}\log f(y_{i}-y_{j};\theta).\end{split}$
+However, the univariate marginal CL cannot estimate covariance parameters $\theta$ while the pairwise difference CL cannot estimate $\mu$ directly, limiting its impact on PS adjustment.
+In addition, both approaches ignore higher-order spatial dependence, which is essential for accurately approximating Gaussian process models.
+To address these gaps, we explore the application of ISIW to the Vecchia approximation. Let the weights be defined as in (5). We initially considered weighting the likelihood as
 $\begin{split}\log\mathcal{L}_{WV}(\boldsymbol{\psi};\textbf{y})&=w_{p(1)}\log f(y_{p(1)};\boldsymbol{\psi})\\ &+\sum_{i=2}^{n}\left(\prod_{j\in\{p(i)\}\cup q(i)}w_{j}\right)\log f(y_{p(i)},\textbf{y}_{q(i)};\boldsymbol{\psi})\\ &-\sum_{i=2}^{n}\left(\prod_{j\in q(i)}w_{j}\right)\log f(\textbf{y}_{q(i)};\boldsymbol{\psi}),\end{split}$
-to maintain the probabilistic interpretation of the inverse weighting.
-However, numerical issues arise due to taking the product of several small weights, greatly increasing the chance of extreme values.
-As a more computationally efficient approach, we approximate the true weight using the following weighted Vecchia approximation
-$\begin{split}\log\mathcal{L}_{WV}(\boldsymbol{\psi};\textbf{y})&=w_{p(1)}\log f(y_{p(1)};\boldsymbol{\psi})\\ &+\sum_{i=2}^{n}w_{p(i)}\log f(y_{p(i)},\textbf{y}_{q(i)};\boldsymbol{\psi})\\ &-\sum_{i=2}^{n}w_{p(i)}\log f(\textbf{y}_{q(i)};\boldsymbol{\psi}).\end{split}$
-3.3 Numerical estimation
-The ISIW for both the Vecchia and composite likelihood can be estimated by standard optimization procedures.
-In our study, we used the L-BFGS-B routine as implemented in the optim package in the R language.
+to maintain the probabilistic interpretation of the inverse weighting for the component conditional densities.
+However, numerical issues arise due to taking the product of several weights, greatly increasing the chance of extreme values.
+As a more stable approach, we approximate the true weight using
+$\begin{split}\log\mathcal{L}_{WV}(\boldsymbol{\psi};\textbf{y})&=w_{p(1)}\log f(y_{p(1)};\boldsymbol{\psi})\\ &+\sum_{i=2}^{n}w_{p(i)}\log f(y_{p(i)},\textbf{y}_{q(i)};\boldsymbol{\psi})\\ &-\sum_{i=2}^{n}w_{p(i)}\log f(\textbf{y}_{q(i)};\boldsymbol{\psi}).\end{split}$ (6)
+We also considered the pairwise marginal composite likelihood as a candidate for ISIW, as it naturally accommodates weighting while allowing estimation of $\mu$.
+However, preliminary analyses indicated that ISIW applied to the Vecchia approximation outperformed the pairwise marginal CL across all evaluation metrics (Table S2).
+We therefore focus the remainder of the paper on ISIW combined with Vecchia.
+The ISIW parameters can be numerically estimated by standard optimization procedures.
+We used the L-BFGS-B routine as implemented in the optim package in the R language.
+In our experience, other derivative-free optimization procedures such as the Nelder-Mead algorithm also work well.
 Initial values for parameters to be estimated were selected with general rules of thumb from the GPVecchia R package and all other optimization parameters were set to their default settings.
-3.4 Winsorization of extreme weights
-As with other inverse weighting procedures, a key concern for ISIW is extreme weights.
-Observation locations in sparsely sampled areas will have extremely high weights in the final likelihood and can cause numerical issues in the optimization.
-Two straightforward options are trimming and Winsorization.
-Trimming involves eliminating values considered as outliers, while Winsorizing enforces outliers to a threshold value.
-We chose Winsorization to handle extreme values because trimming observations can remove rare information not captured by less extreme weights, weakening the effectiveness of PS adjustment.
-For our Winsorization procedure, we first normalize the sampling intensities $\hat{\lambda}$ to sum to $n$.
-We then Winsorized the normalized intensities by choosing a lower threshold (to be elaborated on later), and normalized the new resultant weights to sum to $n$.
 3.3 Prediction
+Although ISIW involves parameter estimation, the estimated parameters primarily serve as a means for generating probabilistic predictions through the kriging equations.
+It is well known that under standard infill asymptotics, the Gaussian process parameters $\mu$, $\phi$, and $\sigma^{2}$ in a Matérn covariance model are weakly identified and cannot be estimated consistently (zhang_inconsistent_2004; wang_prediction_2020).
+Accordingly, ISIW methods are better suited for improving probabilistic prediction under PS than for recovering the true model parameters.
+Nevertheless an effective CL should aim to estimate as many GP parameters as possible in order to maximize the potential for PS adjustment through weighting.
+In this regard, the Vecchia approximation estimates the full set of mean and covariance parameters, in contrast to univariate marginal and pairwise difference CLs, which is expected to yield improved predictive performance.
+Prediction under ISIW follows by substituting $\hat{\boldsymbol{\psi}}$ obtained by maximizing (6) into the kriging equations in (4).
+This is a key distinction between ISIW and the SLP.
+Whereas prediction by ISIW adheres to kriging by substituting PS-adjusted parameters, the SLP model computes predictions from the estimated distribution of $S\mid\textbf{X, Y}]$, making it much more computationally intensive.
+Prediction variance for ISIW is estimated using the ordinary kriging variance formula (cressie_statistics_2015), which adds a penalty term to the simple kriging variance in (4) for estimating $\mu$.
+The variance is computed by substituting estimates of $\theta$ from maximizing (6) into
+$\begin{split}\sigma^{2}_{WV}(\mathbf{X}_{0})&=C_{\theta}(\textbf{X}_{0},\textbf{X}_{0})-C_{\theta}(\textbf{X}_{0},\textbf{X}_{n})^{\top}\boldsymbol{\Sigma}_{n}^{-1}(\theta)C_{\theta}(\textbf{X}_{0},\textbf{X}_{n})\\ &+\frac{\bigl(1-\mathbf{1}^{\top}{\Sigma}_{n}({\theta})^{-1}{C}_{{\theta}}(\mathbf{X}_{0},\mathbf{X}_{n})\bigr)^{2}}{\mathbf{1}^{\top}{\Sigma}_{n}({\theta})^{-1}\mathbf{1}}.\end{split}$
+schliep_correcting_2023 propose kriging variance estimators under PS, however their method does not apply to our scenario because they seek to recover the variance had the population been sampled.
+In contrast, our goal is to estimate prediction variance as a quantification of uncertainty, rather than to recover an estimate from a superpopulation.
+4 Simulation analysis
+In this section, we present a series of simulation experiments designed to evaluate the performance of ISIW.
+We first describe the evaluation metrics and preliminary analyses used to select a representative KIE from a set of candidates as well as the weight threshold for the winsorization.
+We then conduct describe various simulations comparing ISIW KIE, ISIW KIE COV, and ISIW with the true sampling intensities (ISIW Known, which serves as an oracle implementation) against the MLE and SLP across a range of random field specifications.
+We assess robustness to model misspecification and characterize conditions under which each method performs well for parameter estimation and prediction.
+4.1 Evaluation metrics
+To evaluate the probabilistic predictions of our method, we used the continuous ranked probability score (CRPS) to compare predictions over a test grid while accounting for prediction variance (gneiting_strictly_2007).
+When the predictive distribution is Gaussian, a closed form formula for the CRPS score is
+$\mathrm{CRPS}=\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\left[z_{i}\left(2\Phi(z_{i})-1\right)+2\phi(z_{i})-\frac{1}{\sqrt{\pi}}\right],\qquad z_{i}=\frac{y_{i}-\hat{y}_{i}}{\sigma_{i}},$
+where $\Phi(\cdot)$ and $\phi(\cdot)$ are the standard normal cumulative distribution function and probability density function, respectively; $\sigma_{i}$ is the square root of the prediction variance at test location $\mathbf{x}_{i}$, $y_{i}$ is the true value, and $\hat{y}_{i}$ is the corresponding point prediction.
+As an alternative measure, we also computed root mean squared prediction error (RMSPE).
+To compare parameter estimates, we evaluated bias and root mean squared error (RMSE) relative to the true values.
+4.2 Preliminary analyses
+We first picked a representative composite likelihood and KIE to use in the main simulation study.
+Specifically, we considered the pairwise marginal CL and Vecchia approximation, combined with several KIE choices listed in Table S1.
+Based on a preliminary analysis ranking all combinations by RMSPE over a variety of random field scenarios (Supplementary Section B.1) we chose the Vecchia approximation combined with the CvL.adaptive KIE bandwidth selection criterion (cronie_non-model-based_2018; van_lieshout_infill_2021) which consistently achieved the highest RMSPE ranking (Table S2).
+After selecting an ISIW implementation, we determined the winsorization threshold through a separate preliminary analysis conducted under similar settings to those used in the main simulation experiment (Supplementary Section B.2).
+Based on a sensitivity analysis evaluating predictive performance across candidate quantiles (Tables S5, S6), we winsorized extreme weights to the 93% percentile for all later analyses.
+Winsorization was only applied to weights for ISIW KIE and ISIW KIE COV and not ISIW Known.
+4.3 Point process misspecification
+We generated $B=500$ realizations for two separate Matérn random fields on a $200\times 200$ grid on the unit square: a low-range ($\phi=0.02$) and high-range ($\phi=0.15$) surface with common parameters $\mu=4,\sigma^{2}=1.5,\nu=1,$ and $\tau^{2}=0.1$.
+For each realization, we sampled observation locations according to two distinct point process models with $n=100$: a log Gaussian Cox process (LGCP) and a Thomas process (a parent-offspring point process).
+Traditionally, the SLP has only been evaluated for LGCP designs, where it is correctly specified.
+A small sample size was chosen to align closer with the finite samples encountered in our real data applications.
+For the LGCP, the sampling intensity was given by $\exp\{\beta S(\mathbf{x})\}$.
+For the Thomas process, parent points were first generated from a homogeneous Poisson process with an expected count of $300$.
+The number of offspring for each parent then followed a Poisson distribution with mean $\exp\{\beta S(\mathbf{x})\}$ and offspring points were sampled from a normal distribution centered at their respective parent locations, with a scale parameter of $0.03$.
+The resulting point pattern was conditioned to contain $n=100$ points, just like the LGCP.
+We considered values of $\beta\in\{-1,1,2\}$ to investigate differences under varying strengths of PS.
+For each scenario, we fit the MLE, SLP, ISIW KIE, ISIW KIE COV, and ISIW Known methods.
+We estimated the SLP model parameters using the INLA-SPDE implementation as defined in watson_general_2019 and used penalized complexity (PC) priors on the variance and range parameters (simpson_penalising_2017; fuglstad_constructing_2019).
+We adopted an empirical Bayes approach, specifying priors such that $P(\phi<\phi_{0})=0.50$ and $P(\sigma>\sigma_{0})=0.01$ where $\phi_{0}$ is one quarter of the mean pairwise distance among sampled locations and $\sigma_{0}=\sqrt{0.9\cdot\widehat{\text{Var}}(Y)}$ where $\widehat{\text{Var}}(Y)$ denotes the sample variance of the observed responses.
+The default priors for INLA were used for $\mu$ and $\tau$, set as $N(0,1000)$ and $\text{Gamma}(1,5\times 10^{-5})$, respectively.
+We will refer to the SLP as the INLA-SLP.
+Prediction was evaluated by computing the mean CRPS and RMSPE over the centroids of a $32\times 32$ discretization of the unit square.
+Parameter estimation was assessed by computing bias and RMSE for estimating $\boldsymbol{\psi}:=(\mu,\sigma^{2},\phi,\tau^{2})$.
+For all methods, $\nu$ was assumed to be fixed.
+Prediction results appear in Table 1.
+Across all methods, prediction error increased with stronger PS (larger $|\beta|$) and weaker spatial correlation (smaller $\phi$).
+When the sampling design was correctly specified as a LGCP, INLA-SLP consistently achieved the lowest CRPS across all values of $\beta$ and $\phi$.
+In contrast, under the misspecified Thomas process, ISIW Known yielded the best predictive performance across all scenarios.
+In general, ISIW Known and ISIW KIE predicted better than the MLE, demonstrating the benefit of ISIW for spatial interpolation under PS.
+However, the CRPS for ISIW Known was noticeably lower than that of both ISIW KIE and ISIW KIE COV, indicating a gap between the information contained in the known versus estimated weights.
+In particular, ISIW KIE COV generally underperformed, with CRPS values comparable to the MLE and, in several low range ($\phi=0.02$) scenarios, worse than the MLE.
+The largest gains of ISIW Known compared to INLA-SLP occurred under $\beta=2$ in the misspecified Thomas setting, while ISIW KIE also outperformed INLA-SLP for the Thomas process at $\beta=2$ and performed competitively when $\phi=0.15$ for $\beta=-1$ and $\beta=1$.
+Figure 1 provides an illustrative example comparing the predictive tendencies of MLE, INLA-SLP, and ISIW Known.
+In the data-rich area at the bottom, predictions from all three methods are nearly identical.
+The differences emerge in data-sparse regions where PS is known to have its greatest impact.
+While the MLE overpredicts the center of the grid, the INLA-SLP predicts a much lower value over the same region which decreases predictive error but leads to underestimation in a specific area indicated by the deep purple.
+ISIW, similar to INLA-SLP, lowers predictions in that region but not as drastically, avoiding the central underprediction seen in INLA-SLP but still overestimating in another area, shown as dark orange slightly to the right of the center.
+As shown by this example, both INLA-SLP and ISIW adjust for PS in similar ways, but INLA-SLP tends to apply a larger adjustment in data sparse areas compared to ISIW.
+Table 2 summarizes bias and RMSE for estimation of $\mu$.
+The MLE exhibited substantial bias and RMSE across all settings, with error increasing with larger $|\beta|$ and smaller $\phi$.
+In low range settings, ISIW Known generally achieved the smallest error, whereas INLA-SLP performed best in high range settings across both LGCP and Thomas designs.
+The largest gains of ISIW Known relative to INLA-SLP occurred in low range scenarios.
+ISIW KIE had higher estimation error than INLA-SLP but lower than the MLE, while ISIW KIE COV showed large bias and RMSE overall, consistent with its poor prediction.
+Notably, INLA-SLP had the worst performance under the Thomas process with $\phi=0.02$ and $\beta=2$, coinciding with severe underestimation of $\mu$, underscoring the importance of accurate mean estimation for prediction under PS in finite samples.
+We now summarize results for covariance parameter estimation (Tables S7, S8, S9).
+Overall, trends in covariance parameter estimation were less distinct than those observed for mean estimation, and their relationship with prediction is less clear.
+The MLE was generally the most stable and reliable method for estimating the range and nugget parameters.
+Error for the range was lower when $\phi=0.02$ while error for the nugget was lower when $\phi=0.15$.
+In fact, the nugget was well estimated by all methods when $\phi=0.15$, although the MLE consistently achieved the lowest bias and RMSE.
+For variance estimation, INLA-SLP tended to perform well when $\phi=0.15$ while the MLE performed best when $\phi=0.02$, suggesting a broader trend that INLA-SLP performs better in prediction and inference under stronger spatial correlation.
+However, we found that INLA-SLP was numerical unstable in some low range simulations leading to inflated estimates of covariance parameters, with isolated instances even when $\phi=0.15$.
+None of these numerical issues occurred in the commonly studied scenario of $\beta=1$, $\phi=0.15$ and LGCP design, indicating that the simulation setting traditionally used to evaluate INLA-SLP works quite well.
+ISIW methods generally showed larger error for covariance parameters than both MLE and INLA-SLP and was also numerically unstable in low range scenarios.
+Despite these large estimates for the variance or range, prediction did not suffer too much as evidenced by Table 1.
+These results indicate that low prediction error under PS does not necessarily correspond to accurate covariance parameter estimation compared to accurate estimation of $\mu$.
+4.4 Random field misspecification
+In the previous simulation experiment, we demonstrated how misspecification of the sampling design $\mathbf{X}$ can degrade the performance of INLA-SLP and highlighted the benefit of ISIW, which estimates the sampling intensity nonparametrically.
+We now consider a setting in which the latent random field itself is misspecified, being composed of a sum of multiple random fields representing different components of the underlying process rather than a single Gaussian process with Matérn covariance.
+Because parameter estimation no longer corresponds to the true data-generating mechanism in this setting, we focus exclusively on predictive performance as measured by the CRPS.
+We generated $B=500$ realizations of a latent field $S$ defined as
+$S(\mathbf{x})=S_{\ell}(\mathbf{x})+S_{h}(\mathbf{x})+S_{m}(\mathbf{x}),$
+where $S_{\ell}$, $S_{h}$, and $S_{m}$ are independent Gaussian processes.
+In this construction, $S_{\ell}$ corresponds to a low-range GP, $S_{h}$ to a high-range GP, and $S_{m}$ to a medium-range GP that serves as the shared latent variable governing the sampling intensity.
+Each process has mean zero and marginal variance $\sigma^{2}=0.5$.
+For $S_{\ell}$, we set $\phi=0.02$ and $\nu=1/2$; for $S_{h}$, $\phi=0.30$ and $\nu=2$; and for $S_{m}$, $\phi=0.15$ and $\nu=1$.
+This construction induces heterogeneous smoothness and spatial dependence across scales and generates a latent field that cannot be represented by a single Matérn GP.
+A nugget effect with variance $\tau^{2}=0.05$ was added to each realization of $S$.
+Observation locations were then sampled according to an intensity function of the form $\exp\{\beta S_{m}(\mathbf{x})\}$ with $\beta=1.5$.
+Both LGCP and Thomas point process designs were considered, defined as in the earlier synthetic data experiment, with a fixed sample size of $n=100$.
+Results of the experiment are summarized in Table 3.
+When the sampling design was LGCP, INLA-SLP and ISIW Known achieved nearly identical CRPS values of 0.508 and 0.507, respectively.
+This contrasts with the earlier simulation experiment, in which INLA-SLP consistently outperformed ISIW Known by a substantial margin under the LGCP design.
+Under the misspecified Thomas design, INLA-SLP predicted the worst (CRPS = 0.637), while ISIW methods provide substantial improvements.
+ISIW KIE predicted only marginally better than the MLE, and ISIW KIE COV again underperformed relative to the other ISIW variants.
+4.5 Real data simulation
+To better assess the behavior of ISIW in settings resembling our real data applications, we generated random fields and sampled observations calibrated to match the properties of the Galicia and California AQS datasets, with parameters obtained from INLA-SLP fits to the observed data.
+We first scaled all Easting and Northing coordinates by 100,000 meters.
+For the Galicia data, we sampled $n=63$ points from a SLP with $\mu=2.18$, $\sigma^{2}=0.146$, $\nu=1$, $\phi=0.838$, $\tau^{2}=0.193$, and $\beta=-5.20$ using the same bounding box as the original study region.
+For the California AQS data, we sampled $n=98$ points from a SLP with $\mu=0.88$, $\sigma^{2}=0.567$, $\nu=1$, $\phi=1.15$, $\tau^{2}=0.016$, and $\beta=2.47$.
+Although alternative sampling designs besides the LGCP could be considered, we have already demonstrated that misspecification of either component can affect predictive performance of INLA-SLP.
+The goal of this experiment is therefore to evaluate whether ISIW can still outpredict MLE even when the data are generated from a SLP model derived from the real data.
+Results of the experiment are reported in Table 4.
+Across both regions, all ISIW variants improved predictions relative to the baseline MLE, with ISIW Known achieving the lowest CRPS, followed by ISIW KIE and ISIW KIE COV.
+We noticed larger improvements in prediction error for the ISIW in the Galicia setup compared to the California setup.
+5 Real data analysis
+5.1 Galicia moss biomonitoring
+The Galicia moss dataset has been a widely used example for illustrating the effects of PS on geostatistical inference and prediction (diggle_geostatistical_2010; dinsdale_methods_2019; silva_exact_2024).
+It contains 63 measurements from 1997 and 132 measurements from 2000 of lead concentrations in moss samples, measured in micrograms per gram of dry weight, collected from Galicia, northern Spain (Figure S1).
+Sampling in 1997 was preferential, with a bias toward locations in the north with lower lead concentrations, whereas sampling in 2000 was more regular and non-preferential.
+We fit the MLE, INLA-SLP, ISIW KIE and ISIW KIE COV on the 1997 log-transformed data and generated predictive surfaces over a $35\times 35$ grid covering the Galicia area.
+While previous analyses of the Galicia dataset have used areal models, we elected to use the SPDE approach for INLA-SLP to be consistent with our simulation analysis.
+Figure 2 displays predictive surfaces and corresponding lower and upper quantiles under MLE, INLA-SLP, ISIW KIE, and ISIW KIE COV.
+Across methods, the broad spatial patterns were similar, with lower predicted concentrations in the north and higher levels toward the south and west.
+At a finer scale, MLE and ISIW KIE COV predicted greater spatial heterogeneity driven by the observed data, whereas INLA-SLP and ISIW KIE produced smoother predictive surfaces.
+All three PS methods (INLA-SLP, ISIW KIE, and ISIW KIE COV) were spatially consistent in the direction of adjustment relative to the MLE (Figure 4), differing primarily in the magnitude of correction.
+INLA-SLP produced the largest deviations from the MLE, followed by ISIW KIE and ISIW KIE COV, consistent with the behavior illustrated in the simulation from Figure 1.
+5.2 2010 California Air Quality System (AQS) monitoring
+We estimated average daily $\text{PM}_{2.5}$ concentrations in 2010 using data from 98 EPA AQS monitoring locations in California (Figure S2).
+Figure 3 presents the predictions, with differences relative to the MLE shown in Figure 5.
+All methods once again had similar spatial patterns, with lower concentrations in northern and coastal regions and higher concentrations in the Central Valley and parts of southern California.
+Estimated concentrations generally ranged from approximately 5–10 $\mu\text{g}/\text{m}^{3}$ in northern California to 12–18 $\mu\text{g}/\text{m}^{3}$ in the Central Valley, with higher values in the upper tail of the predictive distribution.
 […]
-4 Simulation analysis
-4.1 Experiment
-We conducted a simulation experiment to evaluate the inferential and predictive performance of ISIW-PM and ISIW-V versus the benchmark MLE and SLP approaches under preferential sampling.
-We generated $B=500$ realizations for two separate Matérn random fields on a $200\times 200$ grid on the unit square: a low-range ($\phi=0.02$) and high-range ($\phi=0.15$) surface with $\mu=4,\sigma^{2}=1.5,\nu=1,$ and $\tau^{2}=0.1$.
-To test robustness of methods against a variety of observation patterns, we sampled three different point processes conditional on having $n=100$ and $n=800$ observations: a log Gaussian Cox process (LGCP), a sigmoidal Cox process (SCP), and a Thomas process.
-Traditionally, the SLP approach has only been tested against LGCP observation patterns where it is correctly specified.
-The intensity function for the LGCP is $\exp\{\beta S(\mathbf{x})\}$.
-For the sigmoidal Cox process, the intensity function follows $\beta(1+\exp\{-S(\mathbf{x})\})^{-1}$.
-For the Thomas process, a set of parent points was first generated from a homogeneous Poisson process.
-The number of offspring for each parent then followed a Poisson distribution with mean $\exp\{\beta S(\mathbf{x})\}$ and offspring points were sampled from a normal distribution centered at their respective parent locations, with a scale parameter of 0.1.
-We set $\beta=1$ for all point processes.
-The SCP, LGCP, and Thomas process were deliberately selected to represent visually similar clustered point processes, ordered by increasing clustering intensity, with more distinct and separated clusters in the Thomas process (Figure 1).
-Figure 1: One realization of the point processes used in the simulation analysis for $n=200$. The top and bottom row fields are low range ($\phi=0.02$) and high range ($\phi=0.15$) respectively.
-For each set of observations, we fit the MLE, SLP, ISIW-PM, and ISIW-V models and compared the results in terms of parameter estimation and prediction.
-The MLE, ISIW-PM, and ISIW-V were estimated using optim as described in Section 3.3.
-We estimated the SLP model parameters using the INLA-SPDE implementation as defined in Watson et al., (2019) and used penalized complexity (PC) priors on the variance and range parameters (Simpson et al.,, 2017; Fuglstad et al.,, 2019).
-For the range parameter we set $P(\phi<0.01)=0.01$ and for the variance parameter we set $P(\sigma>10)=0.01$.
-The default priors for INLA were used for $\mu$ and $\tau$, set as $N(0,1000)$ and $\text{Gamma}(1,5\times 10^{-5})$, respectively.
-Prediction for the MLE and all ISIW methods followed the kriging equations, with point estimates of the requisite parameters substituted into (6).
-We evaluated several variants of the ISIW methods, distinguished by their first-stage sampling intensity weight estimation (Table 1).
-As a gold standard for assessing the optimal performance of ISIW, we also used the true weights extracted from the simulation surface, referred to as “Known” weights.
-The intensity estimators included multiple nonparametric kernel density estimators from spatstat and a parametric LGCP model, which used the same priors and mesh as the SLP.
-To determine a threshold for Winsorization, we conducted a small numerical experiment, simulating a Matérn random field with high range ($\phi=0.15$) 100 times.
-For each replicate, we generated two LGCP point patterns with $\beta=1$, one with $N=100$ and the other with $N=800$.
-ISIW-V was then fit under different weight thresholds using the bw.diggle bandwidth selection, as it was the most prone to numerical instability.
-The best-performing threshold, $1\times 10^{-2}$, was selected based on yielding the best average root mean-squared prediction error (RMSPE).
-All subsequent ISIW methods implemented Winsorization using the $1\times 10^{-2}$ threshold.
-Our approach to evaluating each method prioritized predictive performance due to the aforementioned challenges with standard statistical inference in fixed domain asymptotics and preferential sampling.
-The RMSPE was computed over each center point in a $200\times 200$ grid, yielding 40,000 prediction points per method and simulation scenario.
-Parameter estimation was then assessed by computing relative bias and relative root mean-squared error (RMSE) for the parameter vector $(\mu,\sigma^{2},\phi,\tau^{2},\kappa)$.
-4.2 Results
-Across all simulation scenarios, INLA-SLP and ISIW-V Known were the top two predictors by RMSPE, with similar performance to each other and a clear margin over other methods.
-In general, INLA-SLP had better predictive performance in the high range scenario, whereas ISIW-V Known did so in the low range setting (Tables S1 and S2).
-The one exception was the high range scenario for the $N=100$ Thomas process, where ISIW-V Known had both a lower RMSPE (0.694) and a lower standard deviation (0.15) compared to INLA-SLP (0.749 and 0.27) (Table S3).
-However, there were instances of predictive instability in INLA-SLP, as indicated by the high standard deviation of RMSPE, particularly in the $N=100$ low range SCP and high range Thomas process $N=100$ scenarios where INLA-SLP was misspecified.
-The ISIW methods also tended to run faster than the INLA-SLP at both sample sizes and had similar runtimes to the MLE (Table S4).
-We did not experiment with larger sample sizes than $N=800$ because PS is primarily a concern in small to moderate sample sizes; as sample sizes grow sufficiently large, its effects diminish.
-Table 2: Predictive performance of all sixteen methods based on median rank, mean rank, and percentage of total simulations when RMSPE for the method was lower than that of MLE and SLP.
+6 Discussion
+In this work, we investigated ISIW as a practical approach for adjusting PS in model-based geostatistics.
+By combining ISIW weights estimated by KIE with the Vecchia approximation, we developed a computationally efficient method that improves spatial prediction under PS while avoiding the need to fully specify the likelihood of the observation process $\mathbf{X}$.
+Across a broad set of simulation experiments, ISIW outpredicted the MLE and even the SLP model when the sampling design was misspecified.
+The main finding of this study is that ISIW improves spatial prediction across a wide range of PS scenarios.
+These scenarios include both positive and negative PS effects, alternative sampling designs, increasing strengths of PS, and different latent random field structures, including additive Gaussian processes and processes calibrated to resemble our real data applications.
+In particular, ISIW outpredicted SLP when the design followed a non-LGCP pattern, highlighting the sensitivity of the SLP to point process misspecification.
+In contrast, ISIW relies on a nonparametrically estimated sampling intensity, and was shown to still predict well when the true sampling mechanism deviated from standard SLP assumptions.
+Robustness is key for environmental and ecological applications, where sampling designs are often driven by operational constraints and may contain complex clustering that is difficult to emulate using standard random fields.
+Combining ISIW with the Vecchia approximation has clear advantages over earlier pairwise and univariate CL approaches.
+Whereas CL reduces computation by discarding much of the joint dependence in the data, the Vecchia approximation captures higher order correlations while remaining scalable.
+At the same time, the resulting weighting scheme is less intuitive in the Vecchia likelihood and is sensitive to ordering and tuning decisions such as winsorization thresholds, underscoring the need for more guidance on hyperparameter selection in future work.
+Another key finding from our simulation experiment was the disconnect between parameter estimation accuracy and predictive performance under PS.
+In settings with weak spatial dependence, ISIW achieved low error in estimating $\mu$, consistent with theoretical results for weighted likelihood under independence.
+However, estimation was more challenging in the high range setting, implying this theoretical justification weakens under spatial correlation.
+We did not observe analogous results in ISIW covariance parameter estimation, which was highly biased overall, whereas the MLE consistently achieved the lowest covariance parameter estimation error.
+Even so, ISIW frequently predicted better than the MLE.
+This suggests that in finite samples accurate estimation of covariance parameters is not essential for effective spatial prediction under PS compared to that of $\mu$.
+In addition, kriging-based predictors, while not optimal, can perform well under PS even when the fitted model does not correspond to the true generative model.
+Our work also revealed some limitations of the traditional SLP model.
+While INLA-SLP performed well when the LGCP assumption was correct and spatial dependence was strong, it exhibited numerical instability in covariance parameter estimation under low range conditions and was sensitive to misspecification of the sampling design.
+INLA-SLP also tended to apply more extreme adjustments in data-sparse regions than ISIW, sometimes leading to overcorrection.
+Prior specification and sensitivity analyses remain large standing issues for running complex INLA models (Figure S3).
+These properties underscore the importance of evaluating PS methods under a broader range of data-generating mechanisms than those typically considered in the literature.
+ISIW may also be useful in other settings where a latent GP drives the preferential sampling of locations, such as species distribution modeling (SDM) and ecological studies based on presence–absence or count data.
+In these applications, sampling effort is often spatially structured and correlated with the underlying intensity of the process of interest, leading to biased inference and prediction if ignored.
+While existing approaches in these domains typically rely on binary or count-based models, our work focuses on PS adjustment for continuously observed Gaussian outcomes.
+Nonetheless, the core idea of weighting likelihood contributions to account for PS is broadly applicable, suggesting that extensions of ISIW to SDMs may offer a promising direction for future research.
+The gap between ISIW Known and ISIW with estimated weights indicates that nonparametric intensity estimation remains an area of improvement.
+Our simulations suggest that in finite samples KIEs can recover some but not all of the information needed for optimal adjustment.
+The ISIW KIE COV variant, which incorporates a biased predicted response surface as a spatial covariate in the intensity estimation, did not perform as well as expected.
+While this approach was intentionally exploratory, the results suggest that naïvely incorporating response information into intensity estimation is insufficient.
+More principled methods that integrate response values without requiring full joint likelihood specification may push ISIW KIE performance closer to that of ISIW Known.
+Uncertainty quantification under ISIW remains a challenge.
+We relied on the ordinary kriging variance formula using PS adjusted parameter estimates as a working measure of predictive uncertainty.
+While this approach is pragmatic and likely conservative, and simulations showed that ISIW predicts well compared to the other methods, its coverage properties under PS remain unclear.
+Bootstrap approaches could provide improved uncertainty quantification, but are computationally demanding and require resampling spatial point patterns for which principled procedures are not as well-established.
+Developing methods for ISIW prediction variance is an important direction for future work.
+Despite these limitations, ISIW offers several practical advantages that make it appealing for applied use.
+It is computationally fast, conceptually simple, and easy to integrate with modern Gaussian process approximation methods, many of which already rely on Vecchia-type factorizations.
+Unlike the SLP, ISIW avoids the need to specify and estimate a joint likelihood for the response and sampling locations, reducing both modeling complexity and computational burden.
+Overall, this work suggests that improved spatial prediction under PS can be achieved without full joint modeling of the observation and response processes.
+Further advances in intensity estimation for marked point processes and uncertainty quantification have the potential to enhance ISIW even further for geostatistical analysis under PS.
+Ethics declaration
+Conflict of interest
+The authors have no conflict of interest to declare.
+Figures and Tables
+Table 1: Mean CRPS (SE) across the $B=500$ simulations for the synthetic data experiment with $N=100$ points.
+Bold indicates the best (lowest) value among all methods for the simulation scenario defined in each column.
+Point process Method $\boldsymbol{\beta}=-1$ $\boldsymbol{\beta}=1$ $\boldsymbol{\beta}=2$
+$\phi=0.02$ $\phi=0.15$ $\phi=0.02$ $\phi=0.15$ $\phi=0.02$ $\phi=0.15$
+LGCP MLE 1.035 (0.09) 0.408 (0.09) 1.030 (0.08) 0.409 (0.08) 1.789 (0.15) 0.832 (0.27)
+INLA-SLP 0.687 (0.10) 0.296 (0.04) 0.681 (0.06) 0.296 (0.04) 0.829 (0.13) 0.380 (0.08)
+ISIW Known 0.690 (0.10) 0.352 (0.07) 0.692 (0.10) 0.352 (0.07) 1.086 (0.39) 0.613 (0.26)
+ISIW KIE 0.975 (0.15) 0.386 (0.09) 0.986 (0.17) 0.403 (0.18) 1.689 (0.33) 0.762 (0.30)
+ISIW KIE COV 1.101 (0.14) 0.390 (0.09) 1.123 (0.16) 0.390 (0.08) 2.052 (0.34) 0.806 (0.33)
+Thomas MLE 0.984 (0.09) 0.435 (0.11) 0.981 (0.08) 0.435 (0.11) 1.494 (0.24) 0.840 (0.40)
+INLA-SLP 0.809 (0.31) 0.419 (0.18) 0.728 (0.10) 0.407 (0.14) 1.804 (1.21) 0.786 (0.58)
+ISIW Known 0.688 (0.07) 0.384 (0.08) 0.698 (0.10) 0.384 (0.08) 0.959 (0.31) 0.674 (0.33)
+ISIW KIE 0.913 (0.15) 0.416 (0.10) 0.921 (0.14) 0.416 (0.11) 1.395 (0.28) 0.774 (0.40)
+ISIW KIE COV 1.073 (0.18) 0.420 (0.10) 1.096 (0.19) 0.420 (0.11) 1.737 (0.39) 0.842 (0.42)
+Figure 1: Example predictive surface for the MLE, INLA-SLP, and ISIW KIE approaches for a LGCP point pattern.
+The first panel shows the true field values with observations denoted as points, while the remaining panels display residuals (predicted minus truth) for each method.
+Table 2: Bias (SE) and RMSE (SE) for estimation of $\mu$ across the $B=500$ simulations for the point process misspecification simulation with $n=100$ points.
+For bias, bold indicates the smallest value in absolute magnitude within each point process and simulation scenario; for RMSE, bold indicates the smallest value.
+Metric Point process Method $\boldsymbol{\beta}=-1$ $\boldsymbol{\beta}=1$ $\boldsymbol{\beta}=2$
+$\phi=0.02$ $\phi=0.15$ $\phi=0.02$ $\phi=0.15$ $\phi=0.02$ $\phi=0.15$
+Bias LGCP MLE -1.278 (0.16) -0.451 (0.51) 1.267 (0.16) 0.445 (0.49) 2.229 (0.19) 1.052 (0.57)
+INLA-SLP -0.395 (0.28) -0.017 (0.50) 0.368 (0.27) 0.016 (0.50) 0.604 (0.35) 0.115 (0.56)
+ISIW Known 0.084 (0.41) 0.377 (4.43) -0.088 (0.39) -0.600 (0.65) 0.532 (0.55) 0.067 (0.76)
+ISIW KIE -0.964 (0.45) 0.230 (1.00) 0.891 (0.40) -0.261 (1.23) 1.713 (2.33) 0.534 (0.90)
+ISIW KIE COV -1.204 (0.17) 0.100 (0.61) 1.193 (0.10) -0.105 (0.62) 2.137 (0.20) 0.636 (0.66)
+Thomas MLE -1.174 (0.18) -0.440 (0.50) 1.169 (0.16) 0.437 (0.50) 1.853 (0.29) 0.973 (0.61)
+INLA-SLP 0.378 (0.85) 0.349 (0.41) -0.574 (0.47) -0.339 (0.38) -2.841 (2.01) -0.801 (0.90)
+ISIW Known 0.129 (0.39) 0.599 (0.64) -0.188 (0.40) -0.517 (1.24) 0.457 (0.63) 0.016 (0.80)
+ISIW KIE -0.903 (0.28) 0.148 (0.83) 0.862 (0.46) -0.174 (1.06) 1.698 (0.93) 0.583 (0.78)
+ISIW KIE COV -1.094 (0.19) 0.087 (0.61) 1.090 (0.19) -0.109 (0.64) 1.719 (0.05) 0.050 (9.48)
+RMSE LGCP MLE 1.278 (0.16) 0.555 (0.40) 1.267 (0.16) 0.546 (0.38) 2.229 (0.19) 1.061 (0.55)
+INLA-SLP 0.412 (0.26) 0.410 (0.29) 0.388 (0.24) 0.410 (0.29) 0.623 (0.32) 0.460 (0.34)
+ISIW Known 0.292 (0.30) 0.900 (4.35) 0.288 (0.28) 0.718 (0.52) 0.656 (0.59) 0.595 (0.47)
+ISIW KIE 0.975 (0.43) 0.582 (0.85) 0.915 (0.34) 0.612 (1.10) 1.915 (2.16) 0.743 (0.74)
+ISIW KIE COV 1.204 (0.17) 0.483 (0.38) 1.193 (0.16) 0.489 (0.36) 2.137 (0.20) 0.752 (0.52)
+Thomas MLE 1.174 (0.18) 0.542 (0.39) 1.169 (0.16) 0.536 (0.39) 1.853 (0.29) 0.983 (0.59)
+INLA-SLP 0.717 (0.58) 0.416 (0.34) 0.615 (0.41) 0.405 (0.31) 2.855 (1.99) 0.914 (0.78)
+ISIW Known 0.296 (0.28) 0.717 (0.50) 0.335 (0.31) 0.733 (1.13) 0.630 (0.46) 0.588 (0.54)
+ISIW KIE 0.912 (0.25) 0.564 (0.92) 0.912 (0.52) 0.529 (0.49) 1.698 (0.93) 0.775 (0.78)
+ISIW KIE COV 1.094 (0.19) 0.495 (0.36) 1.090 (0.19) 0.493 (0.37) 1.739 (1.05) 1.205 (9.40)
+Table 3: Mean CRPS (SE) for the multiple Gaussian process simulation experiment.
+Point pattern MLE INLA-SLP ISIW Known ISIW KIE ISIW KIE COV
+LGCP 0.525 (0.06) 0.508 (0.05) 0.507 (0.05) 0.528 (0.10) 0.531 (0.06)
+Thomas 0.540 (0.07) 0.637 (0.18) 0.514 (0.05) 0.530 (0.06) 0.555 (0.08)
+Table 4: Mean CRPS (SE) for the real data simulation.
+Real dataset MLE INLA-SLP ISIW Known ISIW KIE ISIW KIE COV
+Galicia moss 0.212 (0.10) 0.131 (0.09) 0.158 (0.08) 0.171 (0.08) 0.193 (0.10)
+California AQS 0.437 (0.11) 0.261 (0.07) 0.321 (0.13) 0.399 (0.15) 0.415 (0.12)
+Figure 2: Spatial predictions and uncertainty of lead concentrations in Galicia.
+Figure 3: Spatial predictions and uncertainty of air pollution concentrations in California.
+Figure 4: Differences in point predictions between each method and the MLE for the Galicia data. Black points represent observed locations.
+Figure 5: Differences in point predictions between each method and the MLE for the California AQS data. Black points represent observed locations.
+Supplementary Material
+Appendix A Additional details on methods
+A.1 Maximum likelihood estimation
+Justification for geostatistical inference via the MLE derives from spatial large sample theory.
+Two main frameworks have dominated asymptotics for geostatistical estimators: increasing domain asymptotics, and fixed domain (or infill) asymptotics.
+Increasing domain asymptotics assume that as $n\to\infty$, the study region $\mathcal{D}$ expands, ensuring a minimum separation distance between observation locations and a fixed density of observations per unit area.
+Fixed domain asymptotics, on the other hand, keep $\mathcal{D}$ fixed while increasing $n$, leading to a growing observation density and vanishing minimum distance between observations (i.e., observations occur closer together as the sample size increases within the fixed study area).
+Under increasing domain asymptotics, the MLE for $\boldsymbol{\psi}$ is consistent and asymptotically normal (AN) (mardia_maximum_1984; bachoc_asymptotic_2014; bachoc_asymptotic_2020).
+Results are more challenging under fixed domain asymptotics due to the inclusion of increasingly close observations of a spatially correlated process.
+Specifically, for a Matérn Gaussian process with known $\nu$, only a subcomponent of the covariance parameters known as the microergodic parameter is consistently estimable and asymptotically normal (zhang_inconsistent_2004; kaufman_role_2013).
+We refer to this parameter as $\kappa$, given by $\sigma^{2}/\phi^{2\nu}$ for Matérn covariance functions.
+The nugget $\tau^{2}$ is also estimable (tang_identifiability_2021) while the variance $\sigma^{2}$ and range $\phi$ are not (zhang_inconsistent_2004).
+Although the smoothness parameter $\nu$ is theoretically estimable, it is numerically challenging to estimate and is conventionally treated as fixed (loh_fixed-domain_2021).
+Fixed effect coefficients in the mean are typically non-estimable (wang_prediction_2020), except under specific smoothness conditions on the covariates (yu_parametric_2022; bolin_spatial_2024; gilbert_consistency_2025).
+A.2 Composite likelihood
+Let $\textbf{Y}=(Y_{1},Y_{2},...,Y_{n})^{\top}\in\mathbb{R}^{n}$ be a $n\times 1$ random vector with probability density $f(\textbf{y};\boldsymbol{\psi})$ for unknown $r$-dimensional parameter vector $\boldsymbol{\psi}\in\Psi\subseteq\mathbb{R}^{r}$.
+Define $\{\mathcal{A}_{1},...,\mathcal{A}_{K}\}$ to be a set of $K$ marginal or conditional events with associated likelihood $\mathcal{L}_{k}(\theta;\textbf{y})\propto f(\textbf{y}\in\mathcal{A}_{k};\boldsymbol{\psi})$.
+The log composite likelihood (CL) is the weighted sum of the event-specific log likelihoods
+$\log\mathcal{L}_{C}(\boldsymbol{\psi};\textbf{y}):=\sum_{k=1}^{K}w_{k}\log\mathcal{L}_{k}(\boldsymbol{\psi};\textbf{y}),$
+where $\textbf{w}:=(w_{1},...,w_{K})$ is a vector of weights, not necessarily non-negative and the density $f$ follows that defined in (2.1).
+The maximum CL estimate is defined as $\hat{\boldsymbol{\psi}}_{C}:=\text{argmax}_{\boldsymbol{\psi}}\log\mathcal{L}_{C}(\boldsymbol{\psi};\textbf{y})$.
+One common choice of CL is the univariate marginal where the log likelihood is the sum of the log marginal density of each observation.
+$\log\mathcal{L}_{UM}(\boldsymbol{\psi};\textbf{y})=\sum_{i=1}^{n}w_{i}\log f(y_{i};\boldsymbol{\psi}).$
+Typically, the use of this univariate marginal in geostatistics is limited because it ignores dependencies between observations and thus cannot estimate covariance parameters.
+For this reason, pairwise CLs have been much more popular (varin_overview_2011).
+bevilacqua_comparing_2015 compared the efficiency of three different pairwise CL estimators: pairwise marginal (PMLE), pairwise conditional (PCMLE), and pairwise difference (PDMLE) CLs.
+These authors concluded that the PMLE outperformed all other pairwise CLs and recommended its use over the PCMLE and PDMLE.
+The likelihoods for each pairwise CL are
+$\begin{split}\log\mathcal{L}_{PM}(\boldsymbol{\psi};\textbf{y})=\sum_{i<j}w_{ij}\log f(y_{i},y_{j};\boldsymbol{\psi}),\\ \log\mathcal{L}_{PC}(\boldsymbol{\psi};\textbf{y})=\sum_{i\neq j}w_{ij}\log f(y_{i}|y_{j};\boldsymbol{\psi}),\\ \log\mathcal{L}_{PD}(\boldsymbol{\psi};\textbf{y})=\sum_{i<j}w_{ij}\log f(y_{i}-y_{j};\boldsymbol{\psi}).\end{split}$ (7)
+These authors also proved, under increasing domain asymptotics, the consistency and asymptotic normality of $\boldsymbol{\hat{\psi}}_{C}$ estimated from (7).
+bachoc_composite_2019 proved the same properties for the PMLE and PCMLE in the one-dimensional setting under fixed domain asymptotics.
+However, the true utility of CL lies in its computational speed and robustness to misspecification.
+While the MLE requires a $O(n^{3})$ matrix inversion and specification of the joint density, pairwise CL consists only of $O(n^{2})$ terms and only requires correct specification of second order densities.
+The computational efficiency of pairwise CL can be further enhanced by using the weights w to only include pairwise observations within a certain distance $d$ apart (bevilacqua_comparing_2015).
+In this study, we use the PMLE as a basis for ISIW due to its superior performance over other pairwise likelihoods.
+In particular, we do not consider the univariate marginal or PDMLE because they do not directly estimate the entirety of $\boldsymbol{\psi}$.
+A.3 Vecchia approximation
+Theoretical justification of the Vecchia approximation in spatial statistics is relatively light compared to that of the MLE and CL.
+Typical approaches to proving fixed domain asymptotics fail due to the loss of stationarity in the parent process introduced by the Vecchia approximation (zhang_fixed-domain_2024).
+However, it is impossible to deny its excellent empirical performance in applications (heaton_case_2019).
+A key advantage the Vecchia approximation possesses over typical CL is that (2) and (3) correspond to a valid joint probability distribution for Y.
+As a result, the Kullback-Leibler (KL) divergence of the Vecchia approximation with respect to the true distribution can be computed and has been shown to be a nonincreasing function of $m$ when the conditioning sets $q(i)$ are chosen as the nearest neighbors (katzfuss_vecchia_2020; huser_vecchia_2023).
+A.4 Kriging
+Stein established the asymptotic efficiency of kriging under various robustness conditions (stein_asymptotically_1988; stein_uniform_1990; stein_bounds_1990; stein_simple_1993), while wang_prediction_2020 showed that kriging’s prediction error vanishes under a uniform metric.
+Additionally, putter_effect_2001 demonstrated that the difference between predictions using kriging with estimated and the true parameters is asymptotically negligible if the joint Gaussian distributions of the spatial process under the true and estimated covariance functions are contiguous almost surely.
+Given these results, one could argue that Gaussian process models are more appropriate for spatial interpolation rather than inference.
+Appendix B Additional simulation results
+B.1 Choice of kernel estimator
+To select a representative version of the ISIW estimator and improve the interpretability of our simulation, we first conducted a preliminary analysis before the main simulation study.
+The ISIW estimator requires selecting a composite likelihood and a kernel intensity estimator (KIE) for the weights.
+We investigated two likelihoods: the pairwise marginal composite likelihood (PMLE) and Vecchia approximation.
+We denote their ISIW versions by ISIW-PM and ISIW-V respectively.
+We generated $B=500$ Monte Carlo simulations for 12 distinct scenarios based on the SLP model (with fixed parameters $\mu=4$, $\sigma^{2}=1.5$, $\tau^{2}=0.1$, and $\beta=1$).
+These scenarios varied based on:
+  • Sample size of the observation process ($N=100,800$).
+  • Strength of spatial correlation ($\phi=0.02,0.15$).
+  • Parametric form of the underlying point process ($\mathbf{X}$ follows a LGCP, sigmoidal Cox process, or Thomas processs driven by $S$)
+We then ranked each CL-KIE combination based on its RMSPE when predicting the spatial field on a $32\times 32$ grid.
+All considered KIEs and their bandwidth selection strategies are listed in Table S1.
+Based on the RMSPE rankings compiled across these 12 scenarios (Table S2), we determined that the ISIW combination of Vecchia approximation and the CvL.adaptive bandwidth selection yielded the best results.
+Table S1: Point process intensity function estimators considered in the simulation study.
+Method Reference Bandwidth Selection
+diggle diggle_kernel_1985 Least-squares cross-validation
+scott scott_kernel_1992 Rule-of-thumb based on normal reference density
+ppl loader_density_1999 Likelihood cross-validation (leave-one-out)
+CvL cronie_non-model-based_2018 Maximum likelihood cross-validation
+CvL.adaptive van_lieshout_infill_2021 Adaptive bandwidth based on local point density
+R-INLA simpson_going_2016 LGCP model (non-kernel-based)
+Table S2: Predictive performance of all sixteen methods based on median rank, mean rank, and percentage of total simulations when RMSPE for the method was lower than that of MLE and SLP.
 Rank was determined by RMSPE.
@@
 ISIW-PM scott 14 12.7 22.3 24.6
-While ISIW-V using estimated weights did not match the predictive performance of ISIW-V Known, weight estimation variants including diggle and CvL.adaptive showed improved predictive performance over the alternative of no PS adjustment (Table 2).
-They achieved a median rank of 5 out of all 16 methods, had a lower RMSPE than MLE in over 70% of simulations, and even outperformed SLP nearly 30% of the time (for comparison, ISIW-V Known outperformed SLP in 44% of simulations).
-The CvL variant outperformed MLE in 85% of simulations, surpassing even SLP, which did so in 73%.
-Surprisingly, ISIW methods using weights estimated by the parametric LGCP model fitted in INLA performed relatively poorly, with a median rank of 8, though they still consistently outperformed MLE.
-All ISIW-V variants outperformed MLE in over 70% of simulations, whereas ISIW-PM lagged far behind, with some variants on average ranking worse than the MLE.
-The variants that predicted best tended to estimate lower bandwidths which introduced greater noise in prediction, as reflected in the higher standard deviation of RMSPE for these variants.
-In contrast, scott and ppl traded variance for bias, which increased numerical stability but decreased the impact of weighting, making them nearly indistinguishable from prediction based on the MLE.
-Figure 2: Example predictive surface for the MLE, INLA-SLP, and ISIW-V approaches.
-The first panel shows the true field values with observations denoted as points, while the remaining panels display residuals (observed minus predicted values) for each method.
-Figure 2 provides an illustrative example comparing the predictive tendencies of MLE, INLA-SLP, and ISIW-V Known.
-In the data-rich area at the bottom, predictions from all three methods are nearly identical.
-The differences emerge in data-sparse regions where PS is known to have its greatest impact.
-While the MLE overpredicts the center of the grid, the INLA-SLP predicts a much lower value over the same region which decreases predictive error but leads to underestimation in a specific area indicated by the deep purple.
-ISIW-V, similar to INLA-SLP, lowers predictions in that region but not as drastically, avoiding the central underprediction seen in INLA-SLP but still overestimating in another area, shown as dark orange slightly to the right of the center.
-As shown by this example, both INLA-SLP and ISIW-V adjust for PS in similar ways, but INLA-SLP tends to apply a larger adjustment in data sparse areas compared to ISIW-V.
-Table 3: Relative bias and RMSE in parameter estimation for MLE, SLP, and ISIW methods across all simulation scenarios.
+Table S3: Relative bias and RMSE in parameter estimation for MLE, SLP, and ISIW methods across all simulation scenarios for preliminary analysis.
 Bolded values indicate the method with the smallest bias or RMSE for a given parameter.
@@
 ISIW-V Known -0.259 0.409 0.035 20.8 90.9 $4.52\times 10^{3}$ 0.053 1.72 0.712 1.58
-CvL 0.022 0.240 -0.057 1.10 65.1 $2.71\times 10^{3}$0.007 1.14 0.140 0.624
+CvL 0.022 0.240 -0.057 1.10 65.1 $2.71\times 10^{3}$ -0.007 1.14 0.140 0.624
 CvL.adaptive -0.018 0.268 0.362 51.9 $1.25\times 10^{3}$ $3.08\times 10^{4}$ 0.441 2.44 2.44 1.14
@@
 INLA 0.135 0.211 -0.018 0.570 -0.201 0.414 0.859 1.56 0.912 1.69
-Table 3 compares the relative bias and relative RMSE in parameter estimation for the MLE, SLP, and ISIW methods, with a specific focus on the Known, CvL, CvL.adaptive, and INLA variants.
-While the MLE approach overestimated $\mu$ and INLA-SLP estimated $\mu$ accurately, ISIW-V Known significantly underestimated the mean by over 25% (averaging an estimate of three for a true mean of four), with a highly variable RMSE of 0.409.
-The estimated weight variants of ISIW-V had higher error in mean estimation compared to INLA-SLP but lower error than ISIW-V Known and MLE.
-For the range parameter, ISIW-V estimates were consistently too high across all variants and exhibited potential numerical instability.
-In contrast, ISIW-PM underestimated the range but did so with much greater stability.
-MLE had relatively low estimation error for all covariance parameters, particularly $\kappa$ and $\tau^{2}$, aligning with theoretical expectations.
-Aside from the mean, INLA-SLP did not estimate any parameter particularly well, with $\tau^{2}$ showing significant numerical instability.
-There was no clear relationship in the simulation between the method of parameter estimation and resulting predictive performance.
-INLA-SLP estimated the nugget and microergodic parameter extremely poorly, despite these being theoretically among the easiest to estimate, while ISIW-V on average underestimated the mean by 25% and vastly overestimated the range by more than a factor of 90.
-Yet, both methods led in predictive performance.
-In contrast, MLE and the CvL and INLA variants of ISIW produced relatively good parameter estimates but consistently ranked lower in predictive performance, with prediction errors far exceeding those of INLA-SLP and ISIW-V.
-5 Application to the Galicia moss data
-Figure 3: Observed lead concentrations at sampled locations in Galicia in 1997 and 2000.
-The Galicia moss dataset has been a widely used example for illustrating the effects of preferential sampling on geostatistical inference and prediction (Diggle et al.,, 2010; Dinsdale and Salibian-Barrera,, 2019).
-It contains lead concentrations in moss samples, measured in micrograms per gram of dry weight, collected from Galicia, northern Spain.
-Figure 3 shows the spatial distribution of 63 measurements taken in 1997 and 132 measurements taken in 2000.
-Sampling in 1997 was preferential, with a bias toward locations in the north with lower lead concentrations, whereas sampling in 2000 was more regular and non-preferential.
-Further details on the dataset can be found in Diggle et al., (2010).
-We fit each of the MLE, INLA-SLP, ISIW-V, and ISIW-PM approaches using the 1997 and 2000 data separately and generated predictive surfaces over a $20\times 20$ grid covering the Galicia area.
-Based on its good performance in the simulation analysis, we estimated the ISIW first stage intensities using CvL.adaptive bandwidth selection.
-We did encounter numerical difficulties when fitting the INLA-SPDE model to the 1997 data.
-While previous analyses of the Galicia dataset have used areal models, we elected to use the SPDE approach to be consistent with our simulation analysis.
-The model was particularly sensitive to the choice of prior for the range and initial value for the nugget.
-It would sometimes generate predictive surfaces with no variation, where the same value was predicted across the entire region including in locations with observations recording different values.
-Here we present the predictive surface that exhibited some variation and aligned with expectations from kriging, where predictions have similar values to the observations closeby.
-Figure 4: Spatial predictions of lead concentrations in Galicia in 1997 and 2000 using MLE, INLA-SLP, ISIW-V, and ISIW-PM. Points represent the observed data.
-Figure 4 displays the predictions for the lead concentration surface.
-There was little difference for the 2000 data, however, substantial differences emerged in the 1997 data.
-As expected, both the ISIW-V and INLA-SLP estimated higher values in data sparse areas to correct for the preferential sampling of low values.
-The INLA-SLP applied a more extreme correction than the ISIW-V, similar to the example in Figure 2.
-It unexpectedly predicted values exceeding 10.5 in the south, where the nearest observations fall below 5.00.
-It also estimated a very high $\mu$ of 10.79 (Table S5).
-This was unusual considering the maximum observation was 9.51 and the mean of all 63 observations in 1997 was 4.72.
-The MLE, ISIW-V, and ISIW-PM only estimated a $\mu$ of 5.15, 6.07, and 5.95, respectively.
-Besides being a result of INLA-SLP’s PS adjustment, the discrepancy could also be attributed to numerical instability in INLA-SLP estimation, and other priors or initial values might produce different results.
-Similar to our simulation results, the ISIW-PM estimated a lower range compared to other methods.
-This low range likely explains why its predictive surface exhibits little spatial dependence with nearly uniform predictions across the region.
-ISIW-V appears to strike a balance between implementing a moderate adjustment for PS while preserving the spatial correlation evident in the dataset.
-6 Discussion
-In this work, we compared ISIW methods to the SLP and MLE for parameter estimation and prediction under preferential sampling.
-We implemented ISIW in a novel combination with the Vecchia approximation and found that, despite being less well-suited for geostatistical inference, ISIW with the Vecchia approximation nevertheless improves predictive accuracy over the MLE approach and remains competitive with the INLA-SLP approach using the true weights.
-We conducted a comprehensive simulation study that extended beyond previous analyses, evaluating the robustness of ISIW across various settings.
-Its benefits persisted across multiple observation point processes, different levels of spatial dependence in the underlying field, and alternative point process intensity estimators used in the first stage of ISIW.
-We also showed an application to the Galicia moss data where ISIW appears to give more reasonable predictions compared to the standard SLP approach.
-Several factors make ISIW an attractive alternative to SLP for PS adjustment, namely its speed, simplicity, and strong predictive performance.
-Not only does it scale more efficiently with increasing $n$, but in terms of the practical implementation, we found that our version of ISIW-V ran faster than INLA-SLP for combined estimation and prediction, even though we did not implement an optimized version.
-ISIW is also simple to use and understand.
-INLA and TMB require specification of both the prior distributions for hyperparameters and mesh triangulation that can significantly alter results (Righetto et al.,, 2020).
-In contrast, ISIW is a straightforward inverse weighting method familiar to much of the research community that only requires an estimated vector of weights.
-Furthermore, since most effective Gaussian process approximation techniques are already based on the Vecchia approximation (Heaton et al.,, 2019), integrating ISIW into state-of-the-art methods is significantly simpler and more straightforward compared to the SLP.
-These computational and conceptual advantages come without sacrificing predictive performance.
-In our simulation analysis, ISIW achieved improved predictive accuracy over the MLE with estimated weights and performed comparably to the gold standard SLP when the true weights were known.
-In scenarios where the SLP was misspecified, ISIW outperformed it, demonstrating greater robustness and fewer parametric assumptions.
-Many improvements can still be made to ISIW.
-A significant gap remains between the predictive accuracy achieved by ISIW with estimated and true weights.
-The lack of accuracy and theoretical guarantees for the nonparametric intensity estimators used in this study may contribute to their inability to achieve similar performance.
-Moreover, Winsorization is necessary to prevent pathological numerical behavior but remains an undesirable ad hoc step.
-Intensity modelling incorporating the values of the observations rather than just their locations may also lead to improved weight estimation.
-Future work should explore more flexible and robust estimators for point process intensities beyond kernel density estimators and the INLA approach.
-Additionally, ISIW does not yield reliable parameter estimates for the underlying geostatistical model.
-The weighted likelihood adjustment introduces substantial bias and error into parameter estimation, yet unexpectedly improves predictions through kriging.
-The apparent disconnect between parameter estimation and prediction warrants further investigation.
-It remains unclear why parameters obtained from the weighted likelihood adjustment consistently lead to better kriging predictions.
-While kriging is known to lose its predictive optimality under preferential sampling, the results of this study provide insight into conditions under which it can still serve as an effective function approximation tool under relaxed assumptions like PS.
-Finally, we did not address uncertainty quantification for parameter estimates or predictions.
-While the kriging variance formula remains available and could be computed using the estimated parameters, its utility in the presence of preferential sampling is uncertain.
-In addition, geostatistical inference is already challenging and preferential sampling adjustment appears to exacerbate the difficulty.
-As a result, any standard errors for parameter estimates obtained under ISIW are unlikely to have valid coverage properties.
-Future work will focus on developing principled approaches to quantifying uncertainty in both inference and prediction under ISIW.
-7 Supplementary Material
-Table S1: Average RMSPE (standard deviation) for methods under the LGCP simulation. Bolded values indicate the method with the lowest average RMSPE across all simulations for that setting.
-N = 100 N = 800
-Method Weights $\phi=0.02$ $\phi=0.15$ $\phi=0.02$ $\phi=0.15$
-MLE 1.726 (0.11) 0.785 (0.17) 1.306 (0.06) 0.361 (0.07)
-INLA-SLP 1.195 (0.08) 0.553 (0.09) 0.995 (0.04) 0.305 (0.04)
-ISIW-V Known 1.193 (0.11) 0.656 (0.13) 0.908 (0.04) 0.346 (0.06)
-CvL 1.660 (0.12) 0.728 (0.16) 1.238 (0.06) 0.353 (0.06)
-CvL.adaptive 1.639 (0.18) 0.714 (0.20) 1.098 (0.07) 0.362 (0.07)
-INLA 1.686 (0.11) 0.783 (0.17) 1.204 (0.05) 0.355 (0.06)
-diggle 1.616 (0.18) 0.712 (0.19) 1.112 (0.14) 0.355 (0.07)
-ppl 1.664 (0.13) 0.730 (0.23) 1.168 (0.16) 0.361 (0.13)
-scott 1.689 (0.11) 0.737 (0.20) 1.279 (0.06) 0.353 (0.06)
-ISIW-PM Known 1.210 (0.10) 0.773 (0.18) 1.066 (0.06) 0.384 (0.08)
-CvL 1.736 (0.12) 0.846 (0.21) 1.390 (0.07) 0.390 (0.09)
-CvL.adaptive 1.614 (0.16) 0.766 (0.18) 1.175 (0.06) 0.395 (0.09)
-INLA 1.764 (0.12) 1.006 (0.26) 1.331 (0.06) 0.392 (0.09)
-diggle 1.593 (0.16) 0.774 (0.19) 1.136 (0.06) 0.382 (0.08)
-ppl 1.723 (0.13) 0.804 (0.18) 1.238 (0.06) 0.380 (0.08)
-scott 1.788 (0.12) 0.845 (0.19) 1.448 (0.07) 0.393 (0.09)
-Table S2: Average RMSPE (standard deviation) for methods under the SCP simulation. Bolded values indicate the method with the lowest average RMSPE across all simulations for that setting.
-N = 100 N = 800
-Method Weights $\phi=0.02$ $\phi=0.15$ $\phi=0.02$ $\phi=0.15$
-MLE 1.289 (0.06) 0.579 (0.10) 0.999 (0.04) 0.271 (0.03)
-INLA-SLP 1.374 (0.31) 0.521 (0.08) 0.935 (0.03) 0.258 (0.02)
-ISIW-V Known 1.228 (0.08) 0.551 (0.08) 0.889 (0.03) 0.270 (0.03)
-CvL 1.280 (0.07) 0.566 (0.09) 0.975 (0.04) 0.268 (0.03)
-CvL.adaptive 1.292 (0.09) 0.574 (0.09) 0.941 (0.05) 0.272 (0.03)
-INLA 1.289 (0.07) 0.581 (0.10) 0.984 (0.04) 0.270 (0.03)
-diggle 1.286 (0.08) 0.570 (0.10) 0.976 (0.04) 0.268 (0.03)
-ppl 1.288 (0.06) 0.569 (0.09) 0.987 (0.04) 0.269 (0.03)
-scott 1.285 (0.06) 0.568 (0.09) 0.995 (0.04) 0.269 (0.03)
-ISIW-PM Known 1.232 (0.09) 0.596 (0.11) 0.900 (0.03) 0.281 (0.03)
-CvL 1.287 (0.07) 0.612 (0.12) 1.017 (0.04) 0.281 (0.04)
-CvL.adaptive 1.287 (0.10) 0.609 (0.12) 0.964 (0.04) 0.278 (0.03)
-INLA 1.298 (0.07) 0.660 (0.14) 1.020 (0.04) 0.287 (0.04)
-diggle 1.282 (0.08) 0.601 (0.11) 1.017 (0.04) 0.281 (0.04)
-ppl 1.303 (0.06) 0.630 (0.12) 1.033 (0.04) 0.283 (0.04)
-scott 1.296 (0.06) 0.628 (0.12) 1.043 (0.04) 0.285 (0.04)
-Table S3: Average RMSPE (standard deviation) for methods under the Thomas process simulation. Bolded values indicate the method with the lowest average RMSPE across all simulations for that setting.
-N = 100 N = 800
-Method Weights $\phi=0.02$ $\phi=0.15$ $\phi=0.02$ $\phi=0.15$
-MLE 1.684 (0.11) 0.822 (0.20) 1.324 (0.06) 0.376 (0.07)
-INLA-SLP 1.214 (0.11) 0.749 (0.27) 0.983 (0.04) 0.324 (0.07)
-ISIW-V Known 1.207 (0.11) 0.694 (0.15) 0.918 (0.06) 0.359 (0.06)
-CvL 1.641 (0.14) 0.765 (0.18) 1.265 (0.06) 0.367 (0.07)
-CvL.adaptive 1.627 (0.20) 0.752 (0.18) 1.121 (0.09) 0.375 (0.07)
-INLA 1.629 (0.12) 0.812 (0.20) 1.216 (0.06) 0.369 (0.07)
-diggle 1.580 (0.18) 0.766 (0.20) 1.137 (0.14) 0.367 (0.07)
-ppl 1.682 (0.26) 0.909 (0.41) 1.278 (0.28) 0.412 (0.25)
-scott 1.644 (0.13) 0.772 (0.24) 1.294 (0.06) 0.367 (0.07)
-ISIW-PM Known 1.217 (0.08) 0.816 (0.21) 1.067 (0.06) 0.403 (0.09)
-CvL 1.744 (0.15) 0.918 (0.28) 1.435 (0.08) 0.414 (0.11)
-CvL.adaptive 1.603 (0.17) 0.816 (0.21) 1.182 (0.07) 0.409 (0.10)
-INLA 1.720 (0.12) 1.040 (0.28) 1.337 (0.06) 0.411 (0.10)
-diggle 1.558 (0.16) 0.830 (0.22) 1.168 (0.07) 0.397 (0.09)
-ppl 1.617 (0.18) 0.816 (0.21) 1.239 (0.07) 0.397 (0.09)
-scott 1.762 (0.13) 0.886 (0.22) 1.482 (0.08) 0.413 (0.10)
 Table S4: Mean (SD) and Median (IQR) runtime in seconds over all simulations for methods under different sample sizes.
@@
 scott 0.25 (0.05) 0.25 (0.05) 15.0 (2.3) 14.8 (3.3)
+B.2 Sensitivity analysis for winsorization threshold
+Once a representative KIE estimator was chosen, we conducted a sensitivity analysis of the choice of threshold for the winsorization to mitigate any numerical issues caused by extreme weights in the likelihood adjustment.
+Percentiles from 90% to 99% were considered.
+We ran $B=100$ Monte Carlo simulations under the random field specifications outlined in the point process misspecification simulation portion of the main text, restricted to the SLP model with parameters $(\mu,\sigma^{2},\phi,\nu,\tau^{2})=(4,1.5,0.15,1,0.1)$ and $\beta=1$ or $\beta=2$.
+Based on using both RMSPE and CRPS as evaluation metrics, we chose 93% to use for the remainder of the study.
+Table S5: RMSPE (SE) of ISIW KIE Estimator under different winsorization thresholds.
+The NA cell indicates there were simulation runs that did not converge due to numerical instability.
+For the $\beta=2$ scenario three out of 100 failed while for $\beta=1$ scenario two out of 100 failed.
+RMSPE (SE)
+(% Upper Quantile) ${\beta}=2$ ${\beta}=1$
+90% 1.242 (0.364) 0.741 (0.176)
+91% 1.233 (0.365) 0.755 (0.281)
+92% 1.221 (0.364) 0.750 (0.280)
+93% 1.214 (0.369) 0.749 (0.281)
+94% 1.212 (0.375) 0.767 (0.314)
+95% 1.201 (0.380) 0.762 (0.310)
+96% 1.203 (0.397) 0.789 (0.375)
+97% 1.221 (0.473) 0.783 (0.340)
+98% 1.247 (0.434) 0.821 (0.382)
+99% NA NA
+Table S6: CRPS (SE) of ISIW KIE Estimator under different winsorization thresholds. Same format as Table S5.
+CRPS (SE)
+(% Upper Quantile) ${\beta}=2$ ${\beta}=1$
+90% 0.755 (0.278) 0.394 (0.095)
+91% 0.750 (0.278) 0.413 (0.247)
+92% 0.741 (0.277) 0.411 (0.248)
+93% 0.741 (0.286) 0.415 (0.249)
+94% 0.744 (0.304) 0.435 (0.285)
+95% 0.744 (0.314) 0.435 (0.284)
+96% 0.763 (0.345) 0.466 (0.354)
+97% 0.802 (0.445) 0.467 (0.318)
+98% 0.855 (0.421) 0.510 (0.360)
+99% NA NA
+B.3 Simulation experiment
+Table S7: Bias (SE) and RMSE (SE) for estimation of $\sigma^{2}$ across the $B=500$ simulations for the point process misspecification simulation experiment with $N=100$ points.
+For bias, bold indicates the smallest value in absolute magnitude within each point process and simulation scenario; for RMSE, bold indicates the smallest value.
+The NA* indicates some simulations diverged.
+Metric Point process Method $\boldsymbol{\beta}=-1$ $\boldsymbol{\beta}=1$ $\boldsymbol{\beta}=2$
+$\phi=0.02$ $\phi=0.15$ $\phi=0.02$ $\phi=0.15$ $\phi=0.02$ $\phi=0.15$
+Bias LGCP MLE -0.156 (0.28) -0.461 (0.41) -0.163 (0.26) -0.453 (0.43) -0.581 (0.18) -0.846 (0.29)
+INLA-SLP -0.022 (0.38) -0.423 (0.38) NA* -0.423 (0.41) -0.167 (0.39) -0.633 (0.34)
+ISIW Known -0.511 (0.41) -0.511 (6.95) -0.523 (0.40) -0.808 (0.69) -1.097 (0.26) -1.196 (0.17)
+ISIW KIE -0.397 (0.46) -0.822 (0.49) -0.451 (0.47) 1.338 (47.86) -0.649 (2.04) -1.061 (0.30)
+ISIW KIE COV -0.916 (0.24) -0.850 (0.27) -0.968 (0.28) -0.826 (0.34) -1.260 (0.18) -1.235 (0.16)
+Thomas MLE -0.181 (0.24) -0.434 (0.47) -0.155 (0.27) -0.439 (0.46) -0.584 (0.26) -0.746 (0.54)
+INLA-SLP 13.357 (277.27) -0.324 (0.58) 0.506 (0.73) -0.342 (0.49) 6.345 (46.20) -0.380 (0.89)
+ISIW Known -0.524 (0.37) -0.854 (0.26) -0.534 (0.40) -0.785 (1.92) -1.124 (0.42) -1.232 (0.17)
+ISIW KIE -0.290 (0.39) 0.847 (27.45) -0.270 (0.45) -0.648 (2.28) -0.587 (0.34) -0.895 (0.52)
+ISIW KIE COV -1.008 (0.25) -0.851 (0.36) -1.026 (0.27) -0.882 (0.33) -1.276 (0.22) -0.880 (7.03)
+RMSE LGCP MLE 0.249 (0.20) 0.548 (0.28) 0.241 (0.19) 0.546 (0.30) 0.582 (0.18) 0.857 (0.26)
+INLA-SLP 0.268 (0.26) 0.502 (0.27) NA* 0.515 (0.29) 0.339 (0.26) 0.662 (0.28)
+ISIW Known 0.571 (0.32) 1.172 (6.87) 0.573 (0.33) 0.870 (0.61) 1.100 (0.25) 1.196 (0.17)
+ISIW KIE 0.504 (0.34) 0.879 (0.38) 0.537 (0.37) 3.043 (47.78) 0.937 (1.92) 1.074 (0.25)
+ISIW KIE COV 0.916 (0.24) 0.862 (0.23) 0.968 (0.28) 0.859 (0.25) 1.260 (0.18) 1.235 (0.16)
+Thomas MLE 0.245 (0.17) 0.556 (0.32) 0.255 (0.18) 0.561 (0.29) 0.594 (0.24) 0.846 (0.36)
+INLA-SLP 13.632 (277.26) 0.524 (0.41) 0.651 (0.60) 0.513 (0.31) 6.383 (46.20) 0.758 (0.60)
+ISIW Known 0.569 (0.29) 0.862 (0.23) 0.588 (0.32) 0.967 (1.83) 1.155 (0.32) 1.232 (0.17)
+ISIW KIE 0.397 (0.28) 2.357 (27.36) 0.421 (0.32) 0.915 (2.19) 0.624 (0.27) 0.971 (0.35)
+ISIW KIE COV 1.008 (0.25) 0.875 (0.29) 1.026 (0.27) 0.906 (0.26) 1.282 (0.18) 1.551 (6.92)
+Table S8: Bias (SE) and RMSE (SE) for estimation of the range parameter $\phi$ across the $B=500$ simulations for the point process misspecification simulation experiment with $N=100$ points.
+For bias, bold indicates the smallest value in absolute magnitude within each point process and simulation scenario; for RMSE, bold indicates the smallest value.
+Metric Point process Method $\boldsymbol{\beta}=-1$ $\boldsymbol{\beta}=1$ $\boldsymbol{\beta}=2$
+$\phi=0.02$ $\phi=0.15$ $\phi=0.02$ $\phi=0.15$ $\phi=0.02$ $\phi=0.15$
+Bias LGCP MLE 0.001 (0.01) -0.032 (0.04) 0.000 (0.01) -0.032 (0.04) -0.004 (0.01) -0.061 (0.03)
+INLA-SLP NA* -0.026 (0.03) NA* -0.026 (0.03) 0.005 (0.01) -0.026 (0.03)
+ISIW Known 0.008 (0.05) -0.063 (0.09) 5.998 (133.91) -0.065 (0.03) 201.968 (2406.95) 0.068 (2.57)
+ISIW KIE 207.207 (4627.34) -0.048 (0.15) 78.978 (968.22) 131.120 (2709.06) NA* NA*
+ISIW KIE COV NA* -0.060 (0.04) NA* -0.057 (0.04) NA* 95.741 (1887.07)
+Thomas MLE -0.002 (0.01) -0.032 (0.04) -0.001 (0.01) -0.029 (0.04) -0.008 (0.00) -0.056 (0.06)
+INLA-SLP 0.045 (0.64) -0.046 (0.02) 0.010 (0.01) -0.044 (0.03) 0.028 (0.01) -0.050 (0.02)
+ISIW Known 0.002 (0.04) -0.066 (0.03) 2.261 (41.64) -0.062 (0.05) 18.463 (182.27) 9.139 (205.42)
+ISIW KIE 2.029 (45.22) 0.044 (1.47) 16.706 (226.19) -0.031 (0.20) -0.006 (0.02) 13.117 (212.18)
+ISIW KIE COV NA* -0.059 (0.05) NA* -0.056 (0.05) NA* 13.713 (226.37)
+RMSE LGCP MLE 0.005 (0.01) 0.043 (0.03) 0.004 (0.00) 0.043 (0.03) 0.006 (0.00) 0.064 (0.03)
+INLA-SLP NA* 0.035 (0.02) NA* 0.035 (0.02) 0.007 (0.01) 0.032 (0.02)
+ISIW Known 0.014 (0.04) 0.073 (0.08) 6.004 (133.91) 0.069 (0.03) 201.975 (2406.95) 0.224 (2.56)
+ISIW KIE 207.211 (4627.34) 0.079 (0.14) 78.982 (968.22) 131.239 (2709.05) NA* NA*
+ISIW KIE COV NA* 0.065 (0.03) NA* 0.066 (0.03) NA* 95.910 (1887.06)
+Thomas MLE 0.004 (0.00) 0.043 (0.03) 0.004 (0.00) 0.042 (0.03) 0.009 (0.00) 0.070 (0.05)
+INLA-SLP 0.045 (0.64) 0.047 (0.02) 0.011 (0.01) 0.046 (0.02) 0.028 (0.01) 0.050 (0.02)
+ISIW Known 0.010 (0.03) 0.069 (0.03) 2.268 (41.64) 0.070 (0.04) 18.478 (182.27) 9.307 (205.41)
+ISIW KIE 2.035 (45.22) 0.157 (1.46) 16.711 (226.19) 0.077 (0.19) 0.010 (0.01) 13.259 (212.17)
+ISIW KIE COV NA* 0.067 (0.03) NA* 0.065 (0.03) NA* 13.874 (226.36)
+Table S9: Bias (SE) and RMSE (SE) for estimation of the nugget across the $B=500$ simulations for the point process misspecification simulation experiment with $N=100$ points.
+For bias, bold indicates the smallest value in absolute magnitude within each point process and simulation scenario; for RMSE, bold indicates the smallest value.
+Metric Point process Method $\boldsymbol{\beta}=-1$ $\boldsymbol{\beta}=1$ $\boldsymbol{\beta}=2$
+$\phi=0.02$ $\phi=0.15$ $\phi=0.02$ $\phi=0.15$ $\phi=0.02$ $\phi=0.15$
+Bias LGCP MLE 0.002 (0.15) -0.007 (0.03) -0.004 (0.13) -0.008 (0.03) -0.019 (0.07) -0.008 (0.03)
+INLA-SLP -0.048 (0.49) 7.426 (166.39) NA* -0.016 (0.04) 0.009 (0.51) -0.003 (0.03)
+ISIW Known 0.066 (0.25) -0.015 (0.05) 0.088 (0.30) -0.018 (0.05) 0.069 (0.18) -0.013 (0.05)
+ISIW KIE 0.109 (0.27) -0.018 (0.07) 0.133 (0.31) -0.005 (0.10) 0.091 (0.22) -0.008 (0.05)
+ISIW KIE COV 0.070 (0.20) -0.010 (0.05) 0.105 (0.24) -0.013 (0.04) 0.087 (0.17) -0.012 (0.05)
+Thomas MLE -0.027 (0.09) -0.007 (0.03) -0.027 (0.11) -0.006 (0.03) -0.064 (0.04) -0.011 (0.02)
+INLA-SLP NA* -0.015 (0.03) 237.511 (5311.04) -0.016 (0.03) -0.005 (0.15) 0.003 (0.03)
+ISIW Known 0.026 (0.19) -0.012 (0.04) 0.066 (0.25) -0.010 (0.04) 0.010 (0.14) -0.013 (0.04)
+ISIW KIE 0.018 (0.18) -0.010 (0.05) 0.053 (0.24) -0.008 (0.06) -0.054 (0.06) -0.012 (0.03)
+ISIW KIE COV 0.066 (0.20) -0.009 (0.04) 0.091 (0.23) -0.008 (0.05) 0.027 (0.14) -0.018 (0.05)
+RMSE LGCP MLE 0.108 (0.10) 0.027 (0.02) 0.101 (0.08) 0.026 (0.02) 0.056 (0.04) 0.022 (0.02)
+INLA-SLP 0.127 (0.47) 7.471 (166.39) NA* 0.030 (0.02) 0.089 (0.50) 0.020 (0.02)
+ISIW Known 0.167 (0.19) 0.039 (0.03) 0.185 (0.25) 0.040 (0.03) 0.138 (0.14) 0.042 (0.03)
+ISIW KIE 0.190 (0.22) 0.050 (0.05) 0.214 (0.26) 0.059 (0.07) 0.144 (0.18) 0.039 (0.04)
+ISIW KIE COV 0.151 (0.15) 0.035 (0.03) 0.183 (0.19) 0.035 (0.03) 0.149 (0.12) 0.038 (0.03)
+Thomas MLE 0.082 (0.05) 0.023 (0.02) 0.087 (0.07) 0.023 (0.02) 0.069 (0.03) 0.021 (0.02)
+INLA-SLP NA* 0.030 (0.02) 237.640 (5311.03) 0.028 (0.02) 0.073 (0.13) 0.025 (0.02)
+ISIW Known 0.124 (0.14) 0.032 (0.02) 0.163 (0.21) 0.033 (0.03) 0.094 (0.10) 0.035 (0.03)
+ISIW KIE 0.115 (0.14) 0.038 (0.04) 0.145 (0.20) 0.040 (0.04) 0.069 (0.04) 0.027 (0.02)
+ISIW KIE COV 0.150 (0.14) 0.036 (0.03) 0.173 (0.18) 0.035 (0.03) 0.111 (0.09) 0.043 (0.03)
+Table S10: Mean RMSPE (SE) across the $B=500$ simulations for the point process misspecification simulation experiment with $N=100$ points.
+Bold indicates the best (lowest) value within each point process (LGCP or Thomas) for each simulation scenario.
+Point process Method $\boldsymbol{\beta}=-1$ $\boldsymbol{\beta}=1$ $\boldsymbol{\beta}=2$
+$\phi=0.02$ $\phi=0.15$ $\phi=0.02$ $\phi=0.15$ $\phi=0.02$ $\phi=0.15$
+LGCP MLE 1.731 (0.11) 0.780 (0.18) 1.725 (0.11) 0.782 (0.17) 2.533 (0.15) 1.387 (0.36)
+INLA-SLP 1.212 (0.12) 0.552 (0.08) 1.206 (0.09) 0.553 (0.08) 1.417 (0.16) 0.701 (0.15)
+ISIW Known 1.190 (0.10) 0.660 (0.14) 1.190 (0.11) 0.660 (0.13) 1.598 (0.36) 1.014 (0.33)
+ISIW KIE 1.596 (0.14) 0.718 (0.16) 1.594 (0.15) 0.733 (0.21) 2.325 (0.25) 1.235 (0.37)
+ISIW KIE COV 1.697 (0.12) 0.734 (0.16) 1.697 (0.12) 0.734 (0.15) 2.550 (0.25) 1.260 (0.36)
+Thomas MLE 1.664 (0.12) 0.820 (0.20) 1.661 (0.11) 0.820 (0.20) 2.219 (0.25) 1.374 (0.46)
+INLA-SLP 1.377 (0.36) 0.770 (0.29) 1.292 (0.17) 0.749 (0.24) 2.897 (1.56) 1.281 (0.78)
+ISIW Known 1.197 (0.08) 0.710 (0.15) 1.204 (0.10) 0.709 (0.15) 1.462 (0.31) 1.078 (0.38)
+ISIW KIE 1.540 (0.16) 0.766 (0.18) 1.542 (0.15) 0.763 (0.18) 2.101 (0.29) 1.256 (0.47)
+ISIW KIE COV 1.637 (0.15) 0.773 (0.18) 1.648 (0.10) 0.769 (0.18) 2.234 (0.33) 1.282 (0.53)
+Appendix C Additional real data analysis
+C.1 Galicia moss data
+Figure S1: Observed lead concentrations at sampled locations in Galicia in 1997 and 2000.
 Table S11: Parameter estimates for the Galicia data.
-Year Method $\mu$ $\sigma^{2}$ $\phi$ (km) $\tau^{2}$
-1997 MLE 5.15 2.85 15.35 2.24
-INLA-SLP 10.79 18.19 17.54 0.00007
-ISIW-V 6.07 2.18 8.45 1.04
-ISIW-PM 5.95 1.24 3.99 2.08
-2000 MLE 2.24 1.40 8.90 0.000
-INLA-SLP 2.27 1.36 10.2 0.00013
-ISIW-V 2.24 1.40 8.91 0.000
-ISIW-PM 2.15 1.37 9.47 0.016
+Dataset Method $\mu$ $\sigma^{2}$ $\phi$ $\tau^{2}$ $\beta$
+Galicia MLE 1.546 0.123 0.142 0.108
+ISIW KIE 1.958 0.0418 0.785 0.0619
+ISIW KIE COV 1.581 0.0468 0.189 0.0854
+INLA-SLP 2.164 0.137 0.863 0.192 -5.43
+C.2 California AQS data
+Figure S2: Average daily concentrations of $\text{PM}_{2.5}$ at monitor locations in California in 2010.
+Table S12: Parameter estimates for the California AQS data.
+Dataset Method $\mu$ $\sigma^{2}$ $\phi$ $\tau^{2}$ $\beta$
+California AQS MLE 2.020 0.134 0.472 0.0146
+ISIW KIE 1.805 0.102 0.528 0.0275
+ISIW KIE COV 1.944 0.0597 0.330 0.0276
+INLA-SLP 0.831 0.643 1.260 0.0167 2.43
+C.3 Prior sensitivity analysis for INLA-SLP
+We conducted a small prior sensitivity analysis for the Galicia data to illustrate how spatial predictions are very sensitive to the prior on the variance parameter.
+Reference values of 0.1, 0.5, and 1 were examined in the PC prior, selected based on the Empirical Bayes estimate obtained from the INLA-SLP fit to the Galicia data ($\hat{\sigma}^{2}=0.146$).
+These values were chosen to span both smaller and larger ranges relative to the empirical estimate, thereby assessing the sensitivity of spatial predictions to the prior specification.
+Figure S3: Spatial predictions of log lead concentrations in Galicia in 1997 and 2000 using INLA-SLP for three different PC priors on the variance parameter ($\sigma^{2}$).
+Points represent the observed data.